### v1

# Constrained Contextual Online Decision Making: A Unified Framework

Abstract
Contextual online decision-making problems with constraints appear in a wide range of real-world applications, such as personalized recommendation with resource limits, adaptive experimental design, and decision-making under safety or fairness requirements. In this paper, we investigate a general formulation of sequential decision-making with stage-wise feasibility constraints, where at each round, the learner must select an action based on observed context while ensuring that a problem-specific feasibility criterion is satisfied. We propose a unified algorithmic framework that captures many existing constrained learning problems, including constrained bandits, active learning with label budgets, online hypothesis testing with Type I error control, and model calibration. Central to our approach is the concept of upper counterfactual confidence bounds, which enables the design of practically efficient online algorithms with strong theoretical guarantees using any offline conditional density estimation oracle. Technically, to handle feasibility constraints in complex environments, we introduce a generalized notion of the eluder dimension, extending it from the classical setting based on square loss to a broader class of metric-like probability divergences. This allows us to capture the complexity of various density function classes and characterize the utility regret incurred due to feasibility constraint uncertainty. Our result offers a principled foundation for constrained sequential decision-making in both theory and practice.
1 Introduction
Contextual sequential online experimentation plays a critical role in various domains, including online platforms (Ban et al., 2024), healthcare systems (Zhou et al., 2023), and other industries. In many real-world applications, the decision maker (DM) often operates under some constraints that must be satisfied (Min and Azizan, 2024). For example, there may be safety constraints in power systems (Ceusters et al., 2023), fairness constraints on online platforms (Verma et al., 2024), or risk constraints in portfolio management (Galichet et al., 2013). The presence of such constraints significantly complicates the sequential decision-making process, making it more challenging to manage the inherent exploration-exploitation trade-off.
To address this challenge, for different statistical tasks, a variety of adaptive constrained decision-making algorithms have been proposed, including constrained bandits (Jagerman et al., 2020), active learning (Awasthi et al., 2021), and online hypothesis testing (Malek et al., 2017). Despite the diversity of these applications, they often share a common structural pattern. In each round, the DM observes a context $x_{t}$, such as the symptoms of the patient in a healthcare setting. Based on this information, the DM selects an action $a_{t}$ that is deemed feasible with respect to some estimated constraint. Subsequently, the DM receives feedback in the form of data sampled from the underlying distributions associated with the context-action pair $(x_{t},a_{t})$ and proceeds to the next round.
The core challenges within this general framework stem from two fundamental issues: (i) the uncertainty regarding the true underlying distributions and (ii) the risk of constraint violations due to estimation errors. If the DM had full knowledge of the true distributions, the problem would be reduced to a deterministic optimization problem, which is typically straightforward to solve. However, given only approximate knowledge of the distributions, an action that appears feasible under the estimated model may, in fact, violate constraints when deployed in practice.
In this paper, we introduce a unified framework for general constrained sequential decision-making problems and present an efficient algorithm to handle it. To be specific, we summarize our contributions as follows:
  • A New General Framework: We propose a new framework to handle general online decision-making problems with stage-wise constraints. Beyond the unconstrained model  (Hu et al., 2025), our framework incorporates many sequential decision-making applications, such as constrained bandits, active learning, risk-aware bandits, and online hypothesis testing. All these problems could be addressed by replacing the abstract functionals defined in  Section 3 with specific task-driven functionals.
  • A Novel Conceptual Generalization of Eluder Dimension: Eluder dimension  (Russo and Van Roy, 2013) is a complexity measure often used to characterize the difficulty of optimal exploration in bandits. Moreover, in constrained online learning (Sridharan and Yoo, 2024), it is also used for characterizing the regret caused by conservative exploration. We extend the idea and concept of the eluder dimension  Russo and Van Roy (2013) from the reward function class with squared error to any density model class with any metric-like probability divergence; see  Definition 5.1. This new measure helps us balance aggressive and conservative exploration to handle feasibility.
  • An Efficient Unified Algorithm: We propose a unified algorithm that could easily transfer any offline density estimation oracle into a efficient sequential decision-making algorithm. Unlike in  Hu et al. (2025) where the underlying density class is restricted to a linear model, our algorithm could handle any density model class, including linear and non-linear model classes. This greatly enriches the range of problems that we can solve.
The remainder of this paper is organized as follows. In Section 3, we introduce our general decision-making model and the key conceptual innovation that characterizes the statistical complexity of learning the underlying distributions. In Section 4, we demonstrate how our framework applies to several real-world scenarios, highlighting its broad practical relevance. In Section 7, we present our main theoretical guarantees along with intermediate lemmas that are critical to the proofs. Subsequently, in Section 8 and Section 9, we provide illustrative examples for bounding the generalized eluder dimension and discuss specific instances of density estimation oracles.
2 Related Work
We review related work from both theoretical and application perspectives, presenting them separately.
From a theoretical perspective, our work is closely related to several research areas, including general sequential decision-making, eluder dimension analysis, and online learning under constraints.
Sequential decision-making is a fundamental problem in modern machine learning. Early work such as Bastani and Bayati (2020) focused on sequential decision-making with finite high-dimensional covariates, using contextual bandits as a representative example. From a more general perspective, Foster et al. (2021) and Zhong et al. (2022) investigated the learnability of general model classes by introducing the decision-estimation coefficient and its generalizations to characterize the complexity of interactive decision-making. However, these studies focus primarily on maximizing expected rewards without addressing additional structural constraints.
In contrast, our paper provides a more general framework that accommodates sequential decision-making under stage-wise constraints. Closely related, Hu et al. (2025) proposed an algorithm for unconstrained online decision-making problems based on linear functional regression. Our framework goes further by supporting both linear and nonlinear density classes while also incorporating constraints into the decision process.
Our work generalizes the notion of the eluder dimension, originally introduced by Russo and Van Roy (2013) to study optimal exploration via function approximation in bandit problems. From a theoretical perspective, Li et al. (2022) and Hanneke (2024) provided deeper insights into the eluder dimension by relating it to other complexity measures of function classes, such as the star number and the threshold dimension. On the application side, the eluder dimension has been widely employed in online learning literature, particularly in bandits and reinforcement learning, to characterize generalization error and regret under function approximation (Osband and Van Roy, 2014; Wang et al., 2020b; Jia et al., 2024; Pacchiano, 2024). In this paper, we extend the concept of the eluder dimension beyond reward functions with squared loss to general density models under arbitrary metric-like probability divergences (see Definition 5.1).
The study of online learning with constraints was initiated by Mannor et al. (2009), where the learner aims to maximize expected rewards subject to long-term sample path constraints. Later on, Mahdavi et al. (2012); Castiglioni et al. (2022); Liakopoulos et al. (2019); Yu et al. (2017) studied the regret upper bound and constraint violation in online optimization with long-term constraints. Instead of studying constraint violation, our paper gives a feasibility guarantee such that with high probability, it holds for all rounds. On the other hand, Sridharan and Yoo (2024); Hutchinson et al. (2024) investigated online optimization with per-round constraints. The benchmark in these papers is the best fixed action in hindsight, which is non-contextual and the regret is static, whereas in our paper, we studied the per-context optimal decision-making problem and our benchmark is contextualized dynamic regret. In brief, we have a high probability feasibility guarantee with dynamic regret guarantee in our paper.
Our work is also closely related to the body of research on online learning with general function approximation. Xu and Zeevi (2020) proposed the Upper Counterfactual Confidence Bound (UCCB) algorithm, which constructs upper confidence bounds in the policy space under general reward function approximation. Foster and Rakhlin (2020) addressed the contextual bandit problem using inverse-gap weighting policies, assuming access to an online regression oracle. More recently, Pacchiano (2024) and Jia et al. (2024) examined general contextual bandits with small noise levels and established connections between reward variance and regret. In the context of reinforcement learning, Wang et al. (2020a, b) studied regret minimization under general value function approximation. In contrast, our work focuses on general sequential decision-making with feasibility constraints, leveraging density function approximation oracles.
Another related literature subfield is reinforcement learning under safety constraints. For comprehensive overviews, we refer readers to the surveys by Gu et al. (2022) and Wachi et al. (2024). Wachi and Sui (2020) investigated reinforcement learning with explicit state-wise safety constraints, and Li and Azizan (2024) studied its multi-agent variant and convergence to Nash equilibrium. Furthermore, HasanzadeZonuzy et al. (2021); Amani et al. (2021a, b) analyzed the sample complexity and regret of learning in value function constrained Markov Decision Processes (MDPs). Additionally, Miryoosefi et al. (2019) studied constrained MDPs where policies are required to lie within a convex policy set. Although these works primarily focus on value-based or policy-space constraints under tabular or linear MDPs, our framework is more general and can accommodate a much broader class of feasibility constraints and non-linear model classes.
Furthermore, we remark on some connections and differences between our paper and control theory. Regret Optimal Control (Sabag et al., 2023; Goel and Hassibi, 2021b) is about minimizing the total cost relative to a clairvoyant optimal policy that has foreknowledge of the system dynamics. Recently, Martin et al. (2022) studied this problem under a safety constraint. In the literature on regret-optimal control, the primary focus has been on identifying the optimal controller and characterizing its structural properties (Sabag et al., 2021; Goel and Hassibi, 2021a; Martin et al., 2024). Moreover, optimal control focuses on minimizing cumulative costs generated by a dynamic system along a single trajectory (Li et al., 2019). In our framework, however, contexts do not have system dynamics, so we aim to maximize cumulative rewards through repeated online experimentation, as opposed to minimizing a cumulative cost over a trajectory.
On the application side, our unified algorithm has broad applicability to a variety of problems, including constrained bandits, stream-based active learning, online hypothesis testing, and sequential $L_{1}$ model calibration, among others. In the following, we review existing results related to these representative examples. The details of how these problems can be incorporated into our framework will be presented in Section 4.
Constrained bandits have been extensively studied in both academia and industry, with applications in areas such as online movie recommendation (Balakrishnan et al., ), news recommendation (Li et al., 2010), online pricing (Ao et al., 2025), and market equilibrium computation under incomplete information (Jalota et al., 2023). Various types of constrained stochastic bandits have been explored, depending on the nature of the constraints. For example, knapsack bandits (Badanidiyuru et al., 2018; Wu et al., 2015; Kumar and Kleinberg, 2022) model scenarios where pulling each arm yields both a reward and a random consumption of a global budget. The objective is to maximize cumulative rewards before the budget is exhausted. Another common formulation involves each arm being associated with both a reward distribution and a cost distribution, where the goal is to maximize cumulative rewards while ensuring that the expected cost at each round stays below a specified threshold—this is known as stage-wise constraints, which is also the focus of this paper. Several studies have addressed linear bandits with stage-wise constraints using various approaches, including explore-exploit (Amani et al., 2019), UCB (Pacchiano et al., 2021), and Thompson Sampling (Moradipari et al., 2021). In addition, Wang et al. (2022) investigated pure exploration under constraints in bandit settings. However, prior work has primarily focused on non-contextual problems. In contrast, we study decision-making with arbitrary context distributions and introduce a significantly more general framework that goes beyond optimizing for mean rewards alone.
Active learning focuses on training models with a small subset of labeled data while achieving strong generalization performance. In stream-based active learning, the learner sequentially observes instances drawn from a distribution and must decide in real-time whether to query the label from an oracle. Awasthi et al. (2021) and Ban et al. (2023) studied this setting where labels are generated stochastically, leveraging NTK approximations for neural embeddings. Song et al. (2019) proposed a contextual bandit algorithm with active learning capabilities. For sequential group mean estimation, Aznag et al. (2023) developed active learning algorithms aimed at controlling the mean-square error and variance through strategic sampling. Additionally, Tae et al. (2024) introduced a bandit-based active learning method to improve fairness via selective sampling. We highlight that our framework naturally subsumes stream-based active learning with expected budget constraint as a special case in Section 4.2.
Online Hypothesis Testing concerns making sequential decisions between the null and alternative hypotheses given informative inputs at each round $t$, a problem originally studied by Wald and Wolfowitz (1948) and Robbins (1952). Since then, extensive research has focused on minimizing the expected experiment length for learning near-optimal decision rules (Naghshvar and Javidi, 2013; Chang and Bloch, 2023). Pan et al. (2022) investigated sequential hypothesis testing under type I and type II error constraints, while Lan and Wang (2021) introduced an active learning-style approach by incorporating budget constraints on the number of queries. In contrast to these works, which aim to design optimal policies under varying resource constraints, Malek et al. (2017) studied fixed-horizon multiple hypothesis testing with type I error control. On the application side, sequential hypothesis testing plays a critical role in clinical trials, where it is essential to maximize the probability of correctly diagnosing conditions while controlling the risk of missed diagnoses (Bartroff et al., 2012; Jennison and Turnbull, 1999). Our proposed algorithm is capable of handling sequential hypothesis testing under any fixed horizon setting. See  Section 4.3 for details.
Sequential Model Calibration aims to ensure that a model’s predictions are unbiased, conditional on the predicted value. Roth (2024) provided a comprehensive overview of various calibration error metrics, including average $L_{1}$, $L_{2}$, and $L_{\infty}$ calibration errors, as well as quantile calibration errors (Gopalan et al., 2021; Úrsula Hébert-Johnson et al., 2018). The first sequential calibration algorithm was introduced by Foster and Vohra (1998). Later,  Foster (1999); Úrsula Hébert-Johnson et al. (2018) proposed the notion of multicalibration and developed corresponding algorithms. Additionally, Abernethy et al. (2011) established the equivalence between calibration and Blackwell approachability. Model calibration is also closely related to other areas of uncertainty quantification, such as conformal prediction (Bastani et al., 2022). Within our framework  Section 4.4, the proposed algorithm can be directly applied to sequential model calibration tasks with $L_{1}$ calibration error.
3 Problem Setup
3.1 General Online Decision Making with Constraints
In this section, we formalize our constrained sequential decision-making problem. The decision maker (DM) operates over a context space $\mathcal{X}$ and a finite action space $\mathcal{A}$**We could extend our algorithm to infinite action spaces by invoking the counterfactual action divergences concept. We refer readers to Xu and Zeevi (2020) for details about action divergences. For each state-action pair $(x,a)$, there are two associated random variables with unknown distributions: $f^{*}_{x,a}\in\mathcal{F}$ and $g^{*}_{x,a}\in\mathcal{G}$. Here, $\mathcal{F}$ and $\mathcal{G}$ are known probability density model classes, which we refer to as the utility density class and the constraint density class, respectively.
The interaction proceeds as follows. At each round $t=1,\dots,T$, the DM observes a context $x_{t}$ drawn i.i.d. from an unknown distribution $\mathcal{Q}_{x}$. Based on $x_{t}$, the DM selects a stochastic policy $\pi_{t}(\cdot|x_{t})\in\Delta(\mathcal{A})$ and samples an action $a_{t}$ according to $\pi_{t}$. Subsequently, the DM receives data points $y_{t}\sim f^{*}_{x_{t},a_{t}}$ and $z_{t}\sim g^{*}_{x_{t},a_{t}}$ before moving to the next round.
In general, a stochastic policy at time $t$ is a mapping $\pi_{t}:\mathcal{X}\rightarrow\Delta(\mathcal{A})$. However, since our sequential decision-making problem is per-context, at each round $t$, the DM only needs to specify the distribution $\pi_{t}(\cdot|x_{t})$ for the realized context $x_{t}$, rather than for the entire space $\mathcal{X}$. For notational simplicity, we will sometimes refer to $\pi_{t}(\cdot|x_{t})$ itself as the policy.
The DM’s statistical task involves two functionals, both bounded within $[0,1]$: the utility functional ${\mathcal{T}}_{1}$, defined on $\mathcal{F}$, and the constraint functional ${\mathcal{T}}_{2}$, defined on $\mathcal{G}$. At each round, upon observing context $x$, the DM aims to sample an action that maximizes expected utility while satisfying a constraint.
[…]
$\max_{\pi\in\Delta(\mathcal{A})}\tilde{{\mathcal{T}}_{1}}(f^{*}_{x_{t},\pi})$
subject $\text{subject to}\ \tilde{{\mathcal{T}}_{2}}(g^{*}_{x_{t},\pi})\leq\tau,$
where $\tau$ is some known threshold number.
However, the DM does not have access to the true distributions $f^{*}$ and $g^{*}$ and must instead estimate them from collected data. Due to the inherent stochasticity and estimation errors, it is generally impossible to guarantee feasibility with certainty. Therefore, in this paper, we focus on high-probability feasibility: the DM must ensure that, for some $\delta>0$, with probability at least $1-\delta$, the selected policy sequence $\left\{\pi_{t}\right\}_{t=1}^{T}$ is feasible for all $t\in[T]$. For any natural number $n$, we use $[n]$ to denote the number set $\left\{1,2,\cdots,n\right\}$.
Assuming the DM successfully selects feasible policies with high probability, we evaluate performance using the utility regret, defined as
$\text{Reg}(T)=\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim\mathcal{Q}_{x}}\tilde{{\mathcal{T}}_{1}}(f^{*}_{x_{t},\pi^{*}})-\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim\mathcal{Q}_{x}}\tilde{{\mathcal{T}}_{1}}(f^{*}_{x_{t},\pi}),$
where $\pi^{*}$ denotes the optimal feasible stochastic policy, and $\pi_{t}$ is the policy selected by the DM at round $t$.
This regret quantifies the utility loss incurred by the DM due to uncertainty and exploration, compared to the performance of the optimal feasible policy in hindsight.
Furthermore, to ensure that a feasible action always exists at each round $t$, we assume the presence of a safe action $a_{0}$ for every context $x\in\mathcal{X}$. Specifically, the DM has prior knowledge that this safe action satisfies $\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=c_{0}$ and $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})\leq\tau$ for all $x$. This guarantees that the DM can always select at least one policy that satisfies the constraint, regardless of the uncertainty in estimating $f^{*}$ and $g^{*}$. Such an assumption is standard in constrained decision-making settings to prevent infeasibility arising from distributional uncertainty  (Pacchiano et al., 2021; Kumar and Kleinberg, 2022).
We formally state this as the following safe action assumption:
Assumption 3.1.
For any context $x\in\mathcal{X}$, there exists a fixed action $a_{0}\in\mathcal{A}$ such that
$\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=r_{0},\ \text{and}\ \tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})=c_{0}\leq\tau$
Moreover, the value $c_{0}$ is known to the DM for all $x\in\mathcal{X}$. This safe action provides a fallback option to ensure feasibility at every round, enabling the DM to avoid constraint violations in the absence of reliable estimates of $f^{*}$ and $g^{*}$.
4 Motivating Examples
To better illustrate the scope and relevance of our general decision-making model introduced in Section 3.1, we present several representative real-world applications that naturally fit within our framework. These include constrained bandits, stream-based active learning, sequential hypothesis testing, and online model calibration. Each of these problems, though arising in different domains and driven by distinct practical goals, shares a common structure of making sequential decisions under uncertainty and constraints. We show that, by appropriately specifying the utility and constraint functionals, our framework can unify and generalize existing approaches to these problems, thereby providing a systematic foundation for constrained sequential learning. Detailed formulations of these examples are provided in the remainder of this section.
4.1 Contextual Bandits with Safety Constraints
We consider the contextual bandits with safety constraint setting. At each round $t$, given context $x_{t}$, the agent selects an action $a_{t}\in\mathcal{A}$. The agent receives a reward cost pair $(r_{t},c_{t})$, where $r_{t}=R^{*}(x_{t},a_{t})+\xi_{t}^{r}$ and $c_{t}=C^{*}(x_{t},a_{t})+\xi_{t}^{c}$. Here, $R^{*}$ and $C^{*}$ are true expected reward and cost functions, and $\xi_{t}^{r}$, $\xi_{t}^{c}$ are i.i.d. zero-mean noise terms.
The agent’s objective is to generate a sequence of policies $\pi_{t}$ that maximizes the expected cumulative reward while ensuring the stage-wise constraint
$\mathbb{E}_{a_{t}\sim\pi_{t}.\xi_{t}^{c}}[C^{*}(x_{t},a_{t})+\xi_{t}^{c}]\leq\tau$
holds at each round.
[…]
4.2 Active Learning with Budget Constraint
Active learning aims to maximize information gain about an unknown model under a limited query budget. We consider the stream-based setting, where the decision maker (DM) sequentially decides which data points to label. The following example is a relaxed version of Awasthi et al. (2021).
At each round $t$, a pair $(x_{t},y_{t})$ is drawn i.i.d. from an unknown density $h^{*}\in\mathcal{H}$. Upon observing $x_{t}$, the learning algorithm issues a prediction $a_{t}\in\mathcal{A}$ (possibly stochastic), and simultaneously decides whether to query the label $y_{t}$.
Following Amani et al. (2019), we assume a query cost $c(x,a)$ for each context-prediction pair. We define an augmented action as $(a,i)\in\mathcal{A}\times\left\{0,1\right\}$, where $i=1$ indicates querying the label. The corresponding augmented stochastic policy is $(\pi,p)$, where $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ specifies the prediction distribution and $p:\mathcal{X}\times\mathcal{A}\rightarrow[0,1]$ gives the probability of querying $y$ given covariate prediction pair $(x,a)$. A loss function $l(a,y)$ quantifies prediction error. The utility distribution is given by the posterior label distribution $h^{*}(\cdot|x,a,i)=h^{*}(\cdot|x,a)$, because whether or not querying for the true label does not influence posterior distribution. The utility functional is defined as the negative expected loss:
${\mathcal{T}}_{1}(h^{*}(\cdot|x,a))=-\int_{y\in\mathcal{Y}}l(a,y)h^{*}(y|x,a)dy.$
We impose a per-round expected budget constraint. That is, for each $t\in[T]$, the augmented policy $(\pi_{t},p_{t})$ must satisfy $\sum_{a\in\mathcal{A}}c(x_{t},a)\pi_{t}(a|x_{t})p_{t}(x_{t},a)\leq\tau.$ Here, the constraint distribution for action $(a,i)$ is the delta distribution $\delta(c(x,a)\cdot\mathbb{I}\left(i=1\right))$, where $\mathbb{I}(x)$ stands for the indicator function. The constraint functional is thus:
${\mathcal{T}}_{2}(\delta(c(x,a)\mathbb{I}\left\{i=1\right\}))=c(x,a)\mathbb{I}\left\{i=1\right\},$
and taking the expectation under the augmented policy recovers the constraint expression.
4.3 Online Hypothesis Testing
We consider the online hypothesis testing in healthcare  (Bartroff et al., 2012; Jennison and Turnbull, 1999). A total of $T$ patients arriving sequentially, each associated with symptom $x_{t}$ sampled i.i.d. from $Q_{x}$. The hypothesis classes are =$\mathcal{H}=\left\{H_{0},H_{1}\right\}$, where $H_{0}$ corresponds to being healthy and $H_{1}$ stands for illness, respectively. After receiving $x_{t}$ at every round, the DM selects a diagnostic action $a_{t}\in\left\{0,1\right\}$, where 00 means healthy and $1$ means ill. The true label $w_{t}=\mathbb{I}\left\{x_{t}\in H_{1}\right\}$ is observed only after the action through some other expensive medical methods such as surgery. For any covariate $x$ and action $a$, the associated random variable is a Bernoulli random variable $w(x,a)\sim\text{Ber}(\mathbb{P}(w=a|x))$. We denote the probability mass function of this Bernoulli random variable as $f^{*}_{x,a}$. The DM aims to maximize the statistical power while controlling the level of critical mistakes under some specific level $\alpha$. Formally, the objective is to solve
$\pi_{t}=\mathop{\mathrm{argmax}}_{a_{t}\sim\pi\in\Delta(\left\{0,1\right\})}\mathbb{P}(a_{t}=0|w_{t}=0)$
[…]
Therefore, for any $x,a$, the utility functional ${\mathcal{T}}_{1}$ is ${\mathcal{T}}_{1}(f^{*}_{x,a})=\mathbb{E}\left[\mathbb{I}\left\{w=a=0\right\}|x\right]$. Similarly, the constraint functional is ${\mathcal{T}}_{2}(f^{*}_{x,a})=\mathbb{E}\left[\mathbb{I}\left\{w=1,a=0\right\}|x\right]$.
4.4 Sequential $L_{1}$ Model Calibration
In statistical prediction, calibration requires that the predictions made by an algorithm be accurate when conditioned on the predictions themselves. More specifically, for a target quantile $q\in[0,1]$, there exist various notions of sequential quantile calibration error, including the $L_{1}$, $L_{2}$, and $L_{\infty}$ calibration errors (Roth, 2024).
$L_{1}$ quantile calibration evaluates empirically calibrated predictions $p_{t}$ over sequences of observations and outcomes $(x_{t},y_{t})$ at a given quantile level $q$. We assume the prediction algorithm outputs values on a discrete grid, $p_{t}\in\{0,1/m,2/m,3/m,\cdots,1\}$, which serves as the action space in our general decision-making model.
Consider a transcript $\pi={(p_{1},x_{1},y_{1}),\cdots,(p_{T},x_{T},y_{T})}$ of length $T$. For each $p\in[1/m]$, let $n(\pi,p)=\sum_{t=1}^{T}\mathbb{I}(p_{t}=p)$ denote the number of times the prediction $p_{t}=p$ is made throughout the transcript. Then, the $L_{1}$ average quantile calibration error for target quantile $q\in[0,1]$ is defined as
$Q_{1}(\pi)=\sum_{p\in[1/m]}\frac{n(\pi,p)}{T}\frac{\sum_{t=1}^{T}\mathbb{I}(p_{t}=p)|q-\mathbb{I}(y_{t}\leq p_{t}|)}{n(\pi,p)}=\sum_{t=1}^{T}\frac{1}{T}\sum_{p\in[1/m]}\mathbb{I}(p_{t}=p)|q-\mathbb{I}(y_{t}\leq p_{t})|,$
where $p_{t}$ is a function of $x_{t}$ and the prior history ${(p_{1},x_{1},y_{1}),\cdots,(p_{t-1},x_{t-1},y_{t-1})}$.
For each pair $(x,p)$, the associated probability distribution is the posterior distribution of $\left|q-\mathbb{I}{(y_{t}\leq p_{t})}\right|$. Denoted its pdf as $f(\cdot|x_{t},p_{t})$. The utility functional ${\mathcal{T}}_{1}$ is defined as the expectation of this quantity, i.e., ${\mathcal{T}}_{1}(f(\cdot|x_{t},p_{t}))=\int yf(y|x_{t},p_{t})dy$. and there is no constraint functional in this example.
5 Estimation Oracle &\&& Generalized Eluder Dimension
In this section, we return to our general constrained decision-making model and lay out the conceptual framework for solving it. Our goal is to provide a principled methodology that connects estimation, feasibility, and optimization in a unified structure.
Recall that the utility and constraint functionals, ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, are defined over the true but unknown probability distributions $f^{*}$ and $g^{*}$. Since these distributions are not directly observable, the decision maker must rely on their empirical estimates, denoted by $\widehat{f}$ and $\widehat{g}$, obtained from past data.
A natural and essential requirement in this context is the stability of the functionals under distributional perturbations. Specifically, we assume there exists a suitable measure of distance or divergence between density functions—such as total variation distance, Hellinger divergence, or another metric-like quantity—such that if the estimated distribution $\widehat{f}$ is sufficiently close to the true distribution $f^{*}$, then the corresponding functional values ${\mathcal{T}}_{1}(\widehat{f})$ and ${\mathcal{T}}_{1}(f^{*})$ must also be close. This continuity property is crucial: it ensures that errors in statistical estimation translate to well-controlled errors in the functional outputs, enabling the decision maker to reason about regret and constraint violations in a quantifiable way.
This intuition forms the basis for the design of our algorithm, where accurate estimation, controlled deviation in functional evaluations, and high-probability feasibility guarantees are carefully integrated.
Therefore, we begin by formally defining the notion of “distance” $\mathsf{D}$ between two probability distributions, referred to as a metric-like probability divergence.
Definition 5.1.
A probability divergence $\mathsf{D}$ is metric-like if it is symmetric and satisfies the following
  • $\mathsf{D}(\mathbb{P}||\mathbb{P}^{\prime})\geq 0$ for any two probability distributions $\mathbb{P}$ and $\mathbb{P}^{\prime}$,
  • $\mathsf{D}(\mathbb{P}||\mathbb{P})=0$ for all $\mathbb{P}$,
  • $\mathsf{D}(\mathbb{P}||\mathbb{P}^{\prime\prime})\leq C_{\mathsf{D}}(\mathsf{D}(\mathbb{P}||\mathbb{P}^{\prime})+\mathsf{D}(\mathbb{P}^{\prime}||\mathbb{P}^{\prime\prime}))$ for some absolute value $C_{\mathsf{D}}$.
For example, TV distance $\mathsf{D}_{TV}$; Hellinger distance metric $\mathsf{D}_{H}$; $L^{p}$ distance $\mathsf{D}_{L^{p}}$ all satisfy Definition 5.1. In this paper, we will assume that $\mathsf{D}\leq C$ for some constant $C$.
To relate the error in functional value evaluation to the error in density estimation, we assume that both ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ satisfy a Lipschitz continuity condition with respect to a chosen metric-like probability divergence $\mathsf{D}$.
Assumption 5.2.
For the metric-like divergence $\mathsf{D}$, the functionals ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ are Lipschitz continuous with respect to $\mathsf{D}$, with Lipschitz constant $L_{\mathsf{D}}$. That is,
$|{\mathcal{T}}_{1}(f)-{\mathcal{T}}_{1}(f^{\prime})|\leq L_{\mathsf{D}}\cdot\mathsf{D}(f||f^{\prime}),\ |{\mathcal{T}}_{2}(f)-{\mathcal{T}}_{2}(f^{\prime})|\leq L_{\mathsf{D}}\cdot\mathsf{D}(f||f^{\prime}).$
Intuitively, the assumption 5.2 ensures that small estimation errors in probability distributions measured by the divergence $\mathsf{D}$ transform into proportionally small errors in the evaluations of utility and constraint functionals. This continuity property enables the decision maker to reliably approximate optimal decisions based on estimated distributions.
5.1 Density Estimation Oracle
Assuming that, for the given statistical functionals ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, we have identified a metric-like divergence $\mathsf{D}$ under which both functionals are Lipschitz continuous, our next goal is a density estimation oracle that provides statistical guarantees with respect to $\mathsf{D}$.
We view the estimation oracle as a black-box supervised learning algorithm that outputs estimated densities and is equipped with provable guarantees in terms of the chosen divergence. A formal definition of such an oracle is provided in Definition 5.3.
Definition 5.3.
Given any supervised learning instance $(\mathcal{X}\times\mathcal{A},\mathcal{Y},\mathcal{F})$ and metric-like probability divergence $\mathsf{D}$, an offline probability estimation oracle $\texttt{Alg}_{\mathsf{Off}}=\left\{\texttt{Alg}_{\mathsf{Off}}^{i}\right\}_{i=1}^{n}$ in our setting is a mapping: $(\mathcal{X}\times\mathcal{A},\mathcal{Y})^{t-1}\rightarrow\mathcal{F}$ such that for any sequence $(x_{1},a_{1},y_{1}),\cdots,(x_{N},a_{N},y_{N})$ with $y_{i}\sim f^{*}_{x_{i},a_{i}}$, the sequence of estimators $\widehat{f}^{i}=\texttt{Alg}_{\mathsf{Off}}^{i}(x_{1},a_{1},y_{1},x_{2},a_{2},y_{2},\cdots,x_{i-1},a_{i-1},y_{i-1})$ satisfies
$\sum_{j=1}^{i-1}\mathsf{D}^{2}(\widehat{f}^{i}_{x_{j},a_{j}},f^{*}_{x_{j},a_{j}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},i,\delta),$
with probability at least $1-\delta$. where $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},i,\delta)$ is some number that is only related to probability model class $\mathcal{F}$, sample point number $i$ and $\delta$.
There exists a wide range of offline density estimation oracles in the statistical literature, including methods such as maximum likelihood estimation, least squares estimation, kernel density estimation, and many others. These oracles differ in their assumptions for modeling, estimation strategies, and the types of statistical guarantees they offer. Additional examples and discussion will be provided in Section 9.
In lieu of that, in this work, we assume access to such an offline density estimation oracle, treated as a black-box algorithm, that provides distributional estimates along with theoretical guarantees in terms of the chosen metric-like divergence $\mathsf{D}$. This assumption allows us to focus on the decision-making layer, while relying on the oracle to ensure that estimation errors are appropriately bounded with respect to $\mathsf{D}$.
Assumption 5.4.
We have access to an offline density estimation oracle $\texttt{Alg}_{\mathsf{Off}}$n learning theory, there are two types of oracles: online and offline. Many papers assume access to online oracles with stronger guarantees  (Sridharan and Yoo, 2024; Foster et al., 2021). However, we just assume access to offline density estimation oracle, which makes our result stronger. defined in Definition 5.3 that can be used for estimation in density model classes $\mathcal{F}$ and $\mathcal{G}$. Moreover, we assume that the corresponding values $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},i,\delta)$ and $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{G},i,\delta)$ are known for any $i$ and $\delta$.
5.2 Generalized Eluder Dimension
In any density estimation problem, statistical estimation error is inevitable. This uncertainty significantly complicates the challenge of ensuring feasibility, as feasibility decisions must be made solely on the basis of estimated distributions. To address this issue, we introduce a generalized notion of the eluder dimension, extending the ideas of Russo and Van Roy (2013), to quantify and control the impact of estimation error on constraint satisfaction. Specifically, we begin by formally defining this generalized eluder dimension, followed by two structural lemmas that are of independent theoretical interest.
The original eluder dimension was proposed by Russo and Van Roy (2013) to analyze reward function approximation in bandit and reinforcement learning settings. It captures the inherent difficulty of estimating the mean reward or value function, typically under squared loss. In contrast, our work extends this concept beyond mean estimation under squared loss to general probability distribution estimation under arbitrary metric-like divergences. This generalization broadens the applicability of the eluder dimension from classical least squares regression to a wide class of density estimation problems, enabling its use in constrained decision-making under uncertainty.
Definition 5.5 (Generalized Eluder Dimension).
Let $\mathcal{G}$ be a class of probability distribution models, where given any $g\in\mathcal{G}$, and any context-action pair $(x,a)$, $g_{x,a}$ is a conditional density function. Let $\mathsf{D}(g_{x,a}||g^{\prime}_{x,a})$ denote a metric-like probability divergence between two densities $g$ and $g^{\prime}$.
A context-action pair $(x,a)$ is said to be $\varepsilon$-dependent on a sequence ${(x_{1},a_{1}),\ldots,(x_{n},a_{n})}$ with respect to $\mathsf{D}$ if, for all functions $g,g^{\prime}\in\mathcal{G}$ satisfying
$\sum_{i=1}^{n}\mathsf{D}^{2}(g_{x_{i},a_{i}}||g^{\prime}_{x_{i},a_{i}})\leq\varepsilon^{2},$
it also holds that
$\mathsf{D}^{2}(g_{x,a}||g^{\prime}_{x,a})\leq\varepsilon^{2}$
We say that $(x,a)$ is $\varepsilon$-independent of the sequence ${(x_{1},a_{1}),\ldots,(x_{n},a_{n})}$ if it is not $\varepsilon$-dependent on it.
The $\varepsilon$-generalized-eluder dimension of the class $\mathcal{G}$ with respect to the divergence $\mathsf{D}$, denoted by $\mathrm{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)$, is defined as the length $d$ of the longest sequence of context-action pairs such that, for some $\varepsilon^{\prime}\geq\varepsilon$, each element in the sequence is $\varepsilon^{\prime}$-independent of its predecessors.
Intuitively, a context-action pair $(x,a)$ is said to be $\varepsilon$-independent of a sequence ${(x_{1},a_{1}),\ldots,(x_{n},a_{n})}$ if there exist two functions $g,g^{\prime}\in\mathcal{G}$ that behave similarly on the sequence—i.e., they have small cumulative squared divergence $\sum_{i=1}^{n}\mathsf{D}^{2}(g_{x_{i},a_{i}}||g^{\prime}_{x_{i},a_{i}})$—but still differ significantly at the new pair $(x,a)$, meaning $\mathsf{D}^{2}(g_{x,a}||g^{\prime}_{x,a})$ remains large.
This notion captures the idea that knowledge of the model’s behavior on a set of inputs may not necessarily determine its behavior elsewhere. The eluder dimension, therefore, quantifies the complexity of a model class in terms of how many such $\varepsilon$-independent decisions can arise, and hence reflects the inherent difficulty of exploration under uncertainty in probabilistic settings.
Regarding the generalized eluder dimension, similar to Russo and Van Roy (2013), we have the following lemmas. For better illustration, we first define the amplitude of a probability subclass $\widetilde{C}\subset\mathcal{G}$ at context action pair $(x,a)$ as
$\omega_{\widetilde{C}}(x,a):=\sup_{g,g^{\prime}\in\widetilde{C}}\mathsf{D}(g_{x,a}||g^{\prime}_{x,a}).$
Lemma 5.6.
If $\left\{r_{t}:t\geq 0\right\}$ is a non-decreasing sequence and $\mathcal{G}_{t}=\left\{g\in\mathcal{G}:\sum_{i=1}^{t-1}\mathsf{D}^{2}(g_{x_{i},a_{i}}||\widehat{g}^{t}_{x_{i},a_{i}})\leq r_{t}\right\}$, then with probability $1$, we have
$\sum_{t=1}^{T}\mathbb{I}\left\{\omega_{\mathcal{G}_{t}}(x_{t},a_{t})>\varepsilon\right\}\leq(\frac{4r_{T}^{1/2}}{\varepsilon}+1)\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon).$
Lemma 5.7.
With the same condition as in lemma 5.6, with probability $1$,
$\sum_{t=1}^{T}\omega_{\mathcal{G}_{t}}(x_{t},a_{t})\leq\frac{1}{T}+C\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})r_{T}T}$
These two lemmas essentially establish that the expressive capacity—or ”amplitude”—of any model class can be effectively bounded by its generalized eluder dimension. In other words, the number of statistically distinguishable predictions that a model class can produce under limited data is controlled by this complexity measure. This insight provides a powerful tool for analyzing the learning dynamics of our algorithm under distributional uncertainty. In particular, it will play a central role in the regret analysis presented in Section 7, where we leverage the eluder dimension to quantify the trade-off between exploration and feasibility.
6 Contextual Decision-Making with Constraints
In this section, we present a unified algorithm for sequential decision-making with constraints, assuming access to an offline estimation oracle $\texttt{Alg}_{\mathsf{Off}}$ that provides guarantees in terms of an upper bound $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}$ and the generalized eluder dimension $\mathrm{dim}_{E}(\mathcal{F},\mathsf{D},\varepsilon)$.
To motivate the design of our algorithm, we highlight two fundamental challenges that must be addressed. First, our access is limited to the offline oracle $\texttt{Alg}_{\mathsf{Off}}$, which constrains our ability to perform adaptive estimation of the underlying system during the decision-making process. Second, due to the reliance on estimated models rather than the true distributions, there is an inherent risk that the selected stochastic policies may violate the constraints. Designing an algorithm that overcomes both challenges—limited adaptivity and potential infeasibility—is the core objective of this section.
6.1 Algorithm Design
Several recent works have explored the use of offline oracles for function approximation in contextual bandits and reinforcement learning, including  Simchi-Levi and Xu (2021) and Qian et al. (2024). Building on this line of research, our proposed algorithm follows an upper-confidence-bound (UCB) design principle, adapted to the setting of constrained sequential decision-making.
At a high level, the algorithm operates as follows. At each round $t$, after observing the context $x_{t}$, the decision maker utilizes the estimated densities $\widehat{f}_{t}$ and $\widehat{g}_{t}$—obtained from the offline oracle—to select a stochastic policy that is approximately feasible and maximizes a carefully constructed upper confidence bound. Unlike traditional UCB algorithms, our confidence bound must account for two key sources of uncertainty: (i) the classical exploration-exploitation trade-off and (ii) the potential infeasibility of actions due to estimation error in constraint satisfaction.
Our algorithm draws inspiration from the Upper Counterfactual Confidence Bound (UCCB) framework introduced by Xu and Zeevi (2020), which proposes a principled approach for constructing confidence bounds in contextual bandits with general function classes. Instead of relying solely on observed actions and rewards, UCCB simulates counterfactual trajectories to estimate the value of alternative policies. This idea allows us to construct upper confidence bounds for any stochastic policy $\pi$ at any given context $x$, even without having observed its actual outcome.
Beyond counterfactual simulation, our algorithm incorporates an additional bonus term—based on the generalized eluder dimension—to explicitly control the regret induced by constraint violations. This term captures the statistical complexity of maintaining feasibility under distributional uncertainty and ensures that the algorithm remains robust even when operating with limited information from the offline oracle.
We now present the pseudo-code of Algorithm 1 and explain its components in detail.
Algorithm 1 General Constrained Decision-Making
utility density class $\mathcal{F}$, constrained density class $\mathcal{G}$, round number $T$, context space $\mathcal{X}$, action space $\mathcal{A}$, tuning parameters $\left\{\beta_{t}\right\}$, $\alpha_{r}$, density estimation oracle $\texttt{Alg}_{\mathsf{Off}}$.
for round $t=1,\cdots,K$ do
     Choose action $a_{t}$ regardless of context $x_{t}$.
end for
for round $t=K+1,\cdots$ do
     Compute $\widehat{f}^{t}$ and $\widehat{g}^{t}$ according to the density estimation oracle $\texttt{Alg}_{\mathsf{Off}}$.
     Observe $x_{t}$.
     for $i=K+1,\cdots,t$ do
         Calculate the counterfactual action $\tilde{a}_{t,i}\sim\tilde{\pi}_{t,i}$ iteratively by
[…]
$+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(i,\delta)}\{\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x_{t},\pi})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x_{t},\pi})\}\Bigg{\}}$
     end for
     Sample action $a_{t}\sim\pi_{t}:=\tilde{\pi}_{t,t}$ and observe sample point $y_{t}$ from distribution $f^{*}_{x_{t},a_{t}}$ and $g^{*}_{x_{t},a_{t}}$.
end for
6.2 Algorithm Explanation
First, as a warm-up phase, during the first $K$ rounds, where $K=|\mathcal{A}|$ is the number of actions, we apply each action exactly once, regardless of the observed context $x_{t}$. This ensures that we obtain at least one observation per action, which is essential for the initialization of the density estimation oracle and the construction of valid upper confidence bounds. Although this initial phase may lead to constraint violations (i.e., infeasible actions), such exploration is necessary to ensure adequate statistical coverage of the action space. Without this step, the algorithm would lack the empirical information needed to make reliable utility and feasibility estimates in subsequent rounds.
Observation: Specifying a Single Distribution After Receiving the Context.
In general, a policy $\pi$ in a contextual bandit setting is defined as a mapping from the context space $\mathcal{X}$ to the probability simplex $\Delta(\mathcal{A})$—that is, $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$. Deploying such a policy in full would, in principle, require specifying a distribution $\pi(\cdot|x)$ for every possible context $x\in\mathcal{X}$.
However, in practice, the decision maker receives the context $x_{t}$ at each round $t$ before choosing an action. This means that it is sufficient to specify the distribution $\pi(\cdot|x_{t})$ for the observed context only, rather than for the entire context space. In effect, the learning task at each round reduces to selecting a single probability distribution over actions, conditioned on the realized context.
This observation significantly simplifies the problem and serves as the foundation for the counterfactual upper-confidence-bound (UCB) approach: since we only need to evaluate and compare action distributions at the current context $x_{t}$, we can simulate and construct confidence bounds for hypothetical (counterfactual) policies without requiring full knowledge of the global policy mapping.
Constructing Confidence Bounds via Counterfactual Simulation for Estimated Policies
As introduced in Section 3, the core of our decision-making framework involves solving a per-context optimization problem at each round. However, the underlying context distribution $Q_{x}$ is unknown, and in settings with a continuous context space $\mathcal{X}$ and non-atomic $Q_{x}$, the decision maker (DM) is almost surely unlikely to encounter the same context more than once during the entire interaction process.
To address this, at each round $t>K$, after observing a new context $x_{t}$, we simulate a sequence of counterfactual policy trajectories $\left\{\widetilde{\pi}_{t,i}\right\}_{i=K+1}^{t-1}$ conditioned on $x_{t}$. Intuitively, we ”pretend” that the specific context $x_{t}$ had appeared in all previous rounds and re-simulate the decision-making history using the previously deployed policies. This simulated trajectory over the last $t-K$ rounds allows the DM to construct empirical quantities that resemble what would have been observed had $x_{t}$ been the actual context all along.
Using this simulated trajectory, we can efficiently construct upper confidence bounds for nearly all stochastic policies evaluated at the current context $x_{t}$, despite not having seen $x_{t}$ before.
We now examine the structure of the upper confidence bound in more detail. At round $t$, our selection rule chooses the policy that maximizes the following quantity:
$\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x_{t},\pi})+\beta_{t}\mathbb{E}_{a\sim\pi,\widetilde{a}_{t,j}\sim\widetilde{\pi}_{t,j}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\tilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x_{t},\pi})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x_{t},\pi})\right\}.$
Here:
  • The first term, ${\mathcal{T}}_{1}(\widehat{f}^{t}_{x_{t},\pi})$, represents the estimated utility based on the current density model $\widehat{f}^{t}$.
  • The second term,
    $\beta_{t}\mathbb{E}_{a\sim\pi,\widetilde{a}_{t,j}\sim\widetilde{\pi}_{t,j}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\tilde{a}_{t,j}\right\}+1}\right],$
    acts as a potential-based exploration bonus. It captures the uncertainty due to the limited number of times action $a$ has been ”virtually” selected under the simulated counterfactual history.
  • The third term,
    $\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x_{t},\pi})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x_{t},\pi})\right\},$
    quantifies the uncertainty in constraint satisfaction and penalizes possible infeasibility. It computes the worst-case difference in constraint evaluations over the high-probability confidence set $\widetilde{C}_{\mathcal{G}}(t,\delta)$.
    This selection rule integrates exploration, exploitation, and constraint awareness into a unified scoring mechanism for policy selection.
Handling Feasibility Constraint
It is important to emphasize that feasibility constraints must be respected not only in real trajectories but also in simulated counterfactual ones. Specifically, after observing the context $x_{t}$ at round $t$, the decision maker must restrict the search space from the full stochastic policy set to a subset of estimated feasible policies, and then apply the $\arg\max$ selection rule as specified in Algorithm 1.
To support this, we now elaborate on two key components: the function confidence set $\widetilde{C}_{\mathcal{G}}(t,\delta)$ and the estimated feasible policy set $\widetilde{\Pi}^{t}_{\delta}(x_{t})$.
Recall that our density estimation oracle $\texttt{Alg}_{\mathsf{Off}}$ guarantees, for any datasets $\mathcal{D}_{1}=\left\{(x_{i},a_{i},y_{i})\right\}_{i=1}^{t-1}$ and $\mathcal{D}_{2}=\left\{(x_{i},a_{i},z_{i})\right\}_{i=1}^{t-1}$, the following high-probability bounds hold:
$\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{f}^{t}_{x_{i},a_{i}}||f^{*}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2),$
$\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{g}^{t}_{x_{i},a_{i}}||g^{*}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{G},t,\delta/2).$
Based on this, we define the function confidence sets as follows:
$\widetilde{C}_{\mathcal{G}}(t,\delta):=\left\{g\in\mathcal{G}:\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{g}^{t}_{x_{i},a_{i}}||g_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{G},t,\delta/2t^{3}),\tilde{{\mathcal{T}}_{2}}(g_{x,a_{0}})=c_{0}\right\}.$
$\widetilde{C}_{\mathcal{F}}(t,\delta):=\left\{f\in\mathcal{F}:\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{f}^{t}_{x_{i},a_{i}}||f_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})\right\}.$
Then, by applying a union bound over rounds $t=1,\dots,T$, we can ensure that with probability at least $1-\delta$, the true models $f^{*}\in\widetilde{C}_{\mathcal{F}}(t,\delta)$ and $g^{*}\in\widetilde{C}_{\mathcal{G}}(t,\delta)$ for all $t$.
To define the estimated feasible policy set, we first introduce the following worst-case constraint value function at any context $x$ under a stochastic policy $\pi$:
$\widetilde{V}^{t}_{\delta}(x,\pi):=\max_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi}).$
Then, the estimated feasible policy set is
$\widetilde{\Pi}^{t}_{\delta}(x):=\left\{\pi:\widetilde{V}^{t}_{\delta}(x,\pi)\leq\tau\right\}\cup\left\{a_{0}\right\}.$
This construction ensures that all policies considered by the algorithm satisfy the feasibility constraint with high probability, and that the safe fallback action $a_{0}$ is always available in case no other policy passes the feasibility test.
In Summary, our general decision-making algorithm proceeds as follows. During the initial $K$ rounds—where $K$ is the number of actions—the algorithm performs uniform exploration by selecting each action exactly once. This ensures that every action is observed at least once, which is essential for initializing the density estimation oracle and enabling reliable model estimation.
At each subsequent round $t>K$, the algorithm first computes the updated estimated models $\widehat{f}_{t}$ and $\widehat{g}_{t}$ based on all data collected up to time $t-1$. It then simulates a counterfactual policy trajectory by re-evaluating past policies as if the current context $x_{t}$ had occurred in earlier rounds. This simulation allows the algorithm to estimate how well each policy would have performed at $x_{t}$ under prior deployments.
Using these simulated trajectories, the algorithm constructs an upper confidence bound for each policy under the current context. It then restricts its attention to the set of estimated feasible policies—those that are likely to satisfy the constraint with high probability—and selects the one that maximizes the upper confidence bound.
This design achieves a careful balance between exploration and exploitation, while simultaneously ensuring high-probability feasibility. It leverages the structure of the counterfactual simulation and confidence-based selection to make robust decisions under uncertainty, even in the absence of adaptive online estimation oracle.
7 Theoretical Results
7.1 Statistical Guarantees of the Algorithm
In this section, we present the theoretical guarantees of our proposed algorithm. For clarity of exposition, we first focus on the case where the function classes $\mathcal{F}$ and $\mathcal{G}$ are finite. That is, we assume $|\mathcal{F}|<\infty$ and $|\mathcal{G}|<\infty$. We remark, however, that our results naturally extend to infinite function classes via standard covering number arguments from empirical process theory. This generalization is deferred to Section 7.2.
To analyze the algorithm’s performance, we begin by defining the following random variable. For any $f\in\mathcal{F}$, the random variable $Y_{f,i}:=\left({\mathcal{T}}_{1}(f_{x_{i},a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x_{i},a_{i}})\right)^{2}$ where $x_{i}$ denotes the context observed at round $i$ and $a_{i}$ is the action sampled according to the algorithm’s stochastic policy. Let $\mathcal{H}_{i-1}$ denote the $\sigma$-algebra generated by the interaction history up to round $i-1$, that is, $\mathcal{H}_{i-1}:=\sigma\left(\left\{(x_{j},a_{j},y_{j},z_{j})\right\}_{j=1}^{i-1}\right)$.
The following lemma provides a concentration bound on the average squared functional deviation across previous rounds, which will be instrumental in bounding the regret.
Lemma 7.1.
For any $\delta\in(0,1)$, with probability at least $21-\delta/$ 2,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[\left({\mathcal{T}}_{1}(f_{x_{i},a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x_{i},a_{i}})\right)^{2}|\mathcal{H}_{i-1}\right]\leq 68\log(|\mathcal{F}|t^{3}/\delta)+2\sum_{i=1}^{t-1}Y_{f,i}$
uniformly for all $t$ and $f\in\mathcal{F}$.
This lemma shows that the sum of the conditional expectations of $Y_{f,i}$ can be effectively controlled by a logarithmic term, along with an additional remainder that is typically small in magnitude. This concentration property enables us to bound the cumulative deviation between any candidate model $f$ and the ground truth $f^{*}$ in terms of their functional evaluations. Leveraging this result, we now proceed to establish a bound on the estimation error incurred by our algorithm.
Lemma 7.2.
Consider a measurable contextual bandit algorithm that selects stochastic policy $\pi_{t}$ based on $\mathcal{H}_{t-1}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t>K$ and every stochastic policy $\pi\in\Pi$, the expected estimation error is upper bounded by
$\left|\mathbb{E}_{x\sim Q}\left[\left\{\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi})-\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{68\log(2|\mathcal{F}|t^{3}/\delta)+4L_{\mathsf{D}}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\frac{\delta}{2t^{3}})}.$
We now turn our attention to the feasibility constraint. As a first step, the following lemma establishes that our constructed function confidence set contains the true function with high probability. This result serves as a foundation for analyzing the feasibility guarantees of the algorithm.
Lemma 7.3.
For any $\delta>0$, we define the following function confidence set.
$\tilde{C}_{\mathcal{G}}(t,\delta)=\left\{g\in\mathcal{G}:\sum_{i=1}^{t-1}D^{2}(g_{x_{i},a_{i}}||\widehat{g}^{t}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{G},t,\frac{\delta}{2t^{3}}),{\mathcal{T}}_{2}(g_{x,a_{0}}))=c_{0}\right\},$
where $a_{0}$ is the safe action. Then with probability at least $1-\delta$, $g^{*}\in\tilde{C}_{\mathcal{G}}(t,\delta)$ holds uniformly for all $t$.
From lemma 7.3, we could see that with high probability, the decision policy set $\widetilde{\Pi}_{\delta}^{t}(x)$ selected in Algorithm 1 will be feasible at any $x\in\mathcal{X}$.
Lemma 7.4.
For any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $t$, $\tilde{\Pi}_{\delta}^{t}(x)$ is a feasible policy set.
Proof of lemma 7.4.
This is a direct result of lemma 7.3 by just using the definition of the feasible policy. ∎
The next lemma asserts that, with high probability, the value of our constructed upper confidence bound exceeds that of the true optimal policy. This property is crucial, as it ensures that our policy selection rule—based on maximizing the upper confidence bound—is not only aligned with the classical UCB principle but also remains valid in the presence of feasibility constraints. In other words, even though we restrict our attention to the estimated feasible policy set at each round, the selection mechanism still favors policies with high potential utility while maintaining constraint satisfaction. This lemma thus provides a key justification for the effectiveness of our upper confidence-based selection strategy under uncertainty.
Lemma 7.5.
For any $\delta\in(0,1)$, for any round $t$, denote $x$ as the context in that round, $\pi^{*}$ is the optimal feasible stochastic policy and $\pi_{t}$ is the stochastic policy that our algorithm applies at round $t$. Then, with probability $1-\delta$,
$\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi_{t}})+\beta_{t}\mathbb{E}_{a\sim\pi_{t},\widetilde{a}_{t,j}\sim\widetilde{\pi}_{t,j}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi_{t}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi_{t}})\geq\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}}).$
This lemma plays a critical role in our analysis, as it guarantees that—even when the true optimal feasible policy lies outside the estimated feasible set, our selection rule still yields an optimistic estimate of its expected utility.
Combining all these parts together, we could get our main theorem about the utility regret.
Theorem 7.6.
For any $\delta\in(0,1)$, with probability at least $1-\delta$, setting
$\beta_{t}=\sqrt{\frac{(34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3}))t}{K}},$
we have the following:
  • 1)
    the policy sequence $\left\{\pi_{t}\right\}_{t=1}^{T}$ selected in Algorithm 1 is feasible.
  • 2)
    The expected utility regret of Algorithm 1 is upper bounded by
    $\text{Reg}(T)\leq$ $2\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+2\sum_{t=1}^{T}\frac{K\beta_{t}}{t}$
    $+\alpha_{r}\sum_{t=1}^{T}\mathbb{E}_{x\sim Q}\left[\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}(x,\pi_{t},y))-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}(x,\pi_{t},y))\right]$
    $\leq$ $\mathcal{O}\Bigg{\{}2\beta_{T}(K+\log(T/K))+2K\beta_{T}\log T$
    $+\alpha_{r}\left(\frac{1}{T}+C\min\left\{T,\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}\right)\Bigg{\}}.$
7.2 Extension to Infinite Function Class
In this subsection, we extend our theoretical analysis from the finite density function class case to the more general setting where the model classes $\mathcal{F}$ and $\mathcal{G}$ are infinite. While the finite case allows for simpler analysis through direct enumeration, many practical models—particularly those involving continuous or high-dimensional representations—are naturally infinite in size.
To handle this, for the simplicity of notation, we assume that both $\mathcal{F}$ and $\mathcal{G}$ are parametrized by one set $\Theta\in\mathbb{R}^{d}$, i.e., each model in $\mathcal{F}$ or $\mathcal{G}$ is indexed by a parameter vector from some subset of $\mathbb{R}^{d}$. This parametric structure enables us to control the complexity of the model through its intrinsic dimension, allowing us to apply covering number arguments and uniform convergence tools from empirical process theory.
Assumption 7.7.
The density model classes $\mathcal{F}$ and $\mathcal{G}$ are parameterized by a compact set $\Theta\subset\mathbb{R}^{d}$ and the diameter is bounded by $\Delta$. We assume that $\mathcal{F}$ and $\mathcal{G}$ satisfy
$\mathsf{D}(f_{\theta_{1},x,a}||f_{\theta_{2},x,a})\leq L_{2}||\theta_{1}-\theta_{2}||,$
$\mathsf{D}(g_{\theta_{1},x,a}||g_{\theta_{2},x,a})\leq L_{2}||\theta_{1}-\theta_{2}||,$
uniformly over $x\in\mathcal{X}$ and $a\in\mathcal{A}$.
Under this assumption, we have the following corollary.
Corollary 7.8.
If model classes $\mathcal{F}$ and $\mathcal{G}$ satisfy 7.7. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, setting
$\beta_{t}=\sqrt{\frac{\left\{72\left(d\log(2+\Delta L_{2}t)+\log(\frac{2t^{3}}{\delta})\right)+L_{\mathsf{D}}^{2}+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\frac{\delta}{2t^{3}})\right\}t}{K}},$
we have the following:
  • 1)
    Our selected policy sequence $\left\{\pi_{t}\right\}_{t=1}^{T}$ is feasible for every round $t\in[T]$.
  • 2)
    The expected utility regret of Algorithm 1 is upper bounded by
    $\text{Reg}(T)\leq$ $2\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+2\sum_{t=1}^{T}\frac{K\beta_{t}}{t}$
    $+\alpha_{r}\left[C\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},\frac{1}{T^{2}}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}+1\right]$
    $\leq$ $\mathcal{O}\Bigg{\{}2\beta_{T}(K+\log(T/K))+2K\beta_{T}\log T$
    $+2\alpha_{r}\left(C\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},\frac{1}{T^{2}}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}+1\right)\Bigg{\}}.$
In the following sections, we turn to concrete instantiations of the key components in our framework. First, we provide several illustrative examples of how to bound the generalized eluder dimension under various probability divergences for common density function classes. These results help quantify the statistical complexity of exploration in practical settings. Then, we present several examples of offline density estimation oracles, including those based on regression and maximum likelihood estimation, which can be readily integrated into our algorithmic framework.
8 Bounding Generalized Eluder Dimension
In this section, we provide concrete examples of distribution classes for which the generalized eluder dimension can be effectively bounded. Specifically, we present results for the following settings: (i) linear function classes under the $L_{2}$ distance, (ii) the Gaussian distribution family under the Hellinger distance, and (iii) exponential families also under the Hellinger distance. These examples illustrate how the abstract complexity measure introduced in our framework can be instantiated and controlled in practical model classes.
8.1 Finite Context Space and Action Space
When the context space $\mathcal{X}$ is finite, the $\varepsilon$-generalized eluder dimension with respect to any probability divergence $\mathsf{D}$ is trivially bounded by $|\mathcal{X}|\cdot|\mathcal{A}|$ for any $\varepsilon$. This follows directly from the definition, since the number of distinct context-action pairs is finite and no more than $|\mathcal{X}||\mathcal{A}|$ such pairs can be mutually $\varepsilon$-independent. In other words, once all possible context-action combinations have been exhausted, any new pair must be $\varepsilon$-dependent on the preceding ones. Although this observation is straightforward, it highlights that the complexity of exploration—as captured by the eluder dimension—is inherently limited in finite domains.
8.2 Linear Function Class with $L^{2}$ Distance
[…]
$w_{k}=\sup_{x,a}\left\{\mathsf{D}(f_{x,a}||f^{\prime}_{x,a}):\sqrt{\sum_{i=1}^{k-1}\mathsf{D}^{2}(f_{x_{i},a_{i}}||f^{\prime}_{x_{i},a_{i}})}\leq\varepsilon^{\prime}\right\}>\varepsilon^{\prime}$
Under some regularity assumptions on the basis functions and model class, we can establish the following bound on the generalized eluder dimension for linear density models.
Theorem 8.1.
Suppose we have $f_{x,a}(y)=\theta^{T}\phi_{x,a}(y)$ where $\left\{\phi_{i}\right\}_{i=1}^{d}$ is a candidate density function class with compact support $\Omega$, $\mu(\Omega)=1$. If we have $||\theta||_{2}\leq R$, $||\phi||_{2}\leq S$. Moreover, if there is some $1/2<\alpha\leq 1$ such that for any $(x,a)$, $\exists 1/2\leq\alpha\leq 1$, such that $21\leq\kappa_{2}(m_{x,a})\leq\frac{1}{\alpha}<$ 2. Then, we have
$\text{dim}_{E}(\mathcal{F},\mathsf{D}_{L^{2}},\varepsilon)\leq\frac{3\alpha-2\alpha^{2}}{3\alpha-2\alpha^{2}-1}\frac{e}{e-1}(\log(1+\frac{8R^{2}S^{2}(1/\alpha-1)}{(2-1/\alpha)\varepsilon^{2}})+\log(\frac{3\alpha-2\alpha^{2}}{3\alpha-2\alpha^{2}-1})).$
8.3 Gaussian Distribution Family with Linear Mean under Hellinger Divergence
Next, we consider a Gaussian distribution family of the form
$f_{\theta}(\cdot|x,a)=\mathcal{N}(\theta^{T}\phi_{x,a},\sigma^{2}),$
where $\phi_{x,a}\in\mathbb{R}^{d}$ is a known feature mapping, and $\theta\in\mathbb{R}^{d}$ is a parameter vector, subject to regularity conditions that will be specified later in theorem 8.2. In this setting, we use the Hellinger distance as the probability divergence measure. For reference, the squared Hellinger divergence between two such distributions is given by
$D_{H}^{2}(f_{x,a}||f^{\prime}_{x,a})=1-e^{-\frac{|\phi_{x,a}^{T}(\theta-\theta^{\prime})|^{2}}{8\sigma^{2}}}.$
Using this form, we can derive an upper bound on the generalized eluder dimension $\mathrm{dim}_{E}(\mathcal{F},\mathsf{D}_{H},\varepsilon)$, as stated in theorem 8.2.
Theorem 8.2.
Assuming that $\phi\in\mathbb{R}^{d}$, $||\phi||_{2}\leq 1$, $||\theta||_{2}\leq 1$, and there are known upper and lower bounds on $\underline{\sigma}^{2}\leq\sigma^{2}\leq\overline{\sigma}^{2}$. Then we have,
$\text{dim}_{E}(\mathcal{F},\mathsf{D}_{H},\varepsilon)\leq C\left(\underline{\sigma},d\right)\log(1+\frac{4}{\varepsilon^{2}}).$
8.4 Exponential Family with Hellinger Distance
Finally, we consider a general exponential family model of the form:
$f_{W}^{x,a}=h(y)\exp(\eta_{x,a}^{T}T(y)-A(\eta_{x,a})),$
where the natural parameter is defined as $\eta_{x,a}=W\phi_{x,a}$, with $\phi_{x,a}\in\mathbb{R}^{d}$ representing the feature vector and $W\in\mathbb{R}^{k\times d}$ being a coefficient matrix. Here, $T(y)$ denotes the sufficient statistic of the distribution and $A(\eta_{x,a})$ is the log-partition function.
Under appropriate regularity conditions on the feature map, the sufficient statistic, and the parameter space, we can establish the following bound on the generalized eluder dimension for this model class.
Theorem 8.3.
Assuming that $||W||_{F}\leq\beta$, $||\phi||\leq 1$. $A(\eta)$ function satisfies $\underline{\lambda}I\preceq~{}\nabla^{2}_{\eta}A(\eta)\preceq\overline{\lambda}I$. Then we have,
$\text{dim}_{E}(\mathcal{F},\mathsf{D}_{H},\varepsilon)\leq d\left(\left(1+\sqrt{1/2+\frac{8}{c\underline{\lambda}}}\right)\frac{e}{e-1}\left(\log(1+\frac{2\beta}{1/2(\varepsilon^{\prime})^{2}})\right)+\log\left(1+\sqrt{1/2+\frac{8}{c\underline{\lambda}}}\right)\right).$
9 Density Estimation Oracles and Computation
In this section, we present several examples of offline density estimation oracles along with some computational procedures. These examples illustrate how the estimation error term $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}$—which plays a central role in our unified  Algorithm 1 can be explicitly characterized and bounded in a variety of practical settings.
9.1 Empirical Risk Minimization for Square Loss Regression
We consider a setting where the true conditional density is given by the Gaussian model:
$f_{\theta^{*},x,a}=\mathcal{N}((\theta^{*})^{T}\phi_{x,a},\sigma^{2}),$
where $\theta^{*}\in\Theta\subset\mathbb{R}^{d}$ is an unknown parameter vector, and $\phi_{x,a}\in\mathbb{R}^{d}$ is a known feature map. For simplicity, we assume that the variance $\sigma^{2}>0$ is known.
Given a dataset $\mathcal{D}=\left\{(x_{i},a_{i},y_{i})\right\}_{i=1}^{n}$, we estimate $\theta^{*}$ by solving the standard least squares problem:
$\min_{\theta}\sum_{i=1}^{n}(y_{i}-\theta^{T}\phi_{x_{i},a_{i}})^{2}.$
We denote our estimator as $\widehat{\theta}_{LS}$. Then we have the following theorem.
Theorem 9.1.
The least squares estimation $\widehat{\theta}_{LS}$ satisfies that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{i=1}^{n}\mathsf{D}_{H}^{2}(f^{*}_{x_{i},a_{i}}||\widehat{f}^{n}_{x_{i},a_{i}})\leq\mathcal{O}\left(\sigma^{2}(d+2\sqrt{d\log(1/\delta)}+2\log(1/\delta))\right).$
Thus, we can upper bound the $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},n,\delta)$ by $\left(\sigma^{2}(d+2\sqrt{d\log(1/\delta)}+2\log(1/\delta))\right)$.
9.2 Maximal Likelihood Estimation for Hellinger Distance
In this section, we consider the use of maximum likelihood estimation (MLE) for offline density estimation when the density function class $\mathcal{F}$ is finite. Given a dataset $\mathcal{D}=\left\{(x_{i},a_{i},y_{i})\right\}_{i=1}^{n}$, the MLE procedure selects the density function that maximizes the log-likelihood over the observed data:
$\widehat{f}_{\mathrm{MLE}}=\max_{f\in\mathcal{F}}\sum_{i=1}^{n}\log(f_{x_{i},a_{i}}(y_{i})).$
The following theorem provides a bound on the estimation error associated with $\widehat{f}_{\mathrm{MLE}}$ in terms of the function class complexity and sample size.
Theorem 9.2.
Denote dataset $\mathcal{D}=\left\{(x_{i},a_{i},y_{i})\right\}_{i=1}^{n}$, assuming that the conditional density function class is $\mathcal{F}$, denote the maximum likelihood estimation as $\widehat{f}^{n}$, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\sum_{i=1}^{n}\mathsf{D}_{H}^{2}(\widehat{f}^{n}_{x_{i},a_{i}}||f^{*}_{x_{i},a_{i}})\leq\log(|\mathcal{F}|/\delta).$
Therefore, the $\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta)=\log(|\mathcal{F}|/\delta)$.
To conclude this section, we show that maximum likelihood estimation for exponential family model can be efficiently carried out by solving the associated score equation.
9.3 Score Equation Method for Exponential Families
We now consider the following exponential family model:
$p(y|x,a;\theta)=h(y)\exp\left\{\eta(\theta,x,a)^{T}T(y)-A(\theta,x,a)\right\},$
where $\eta(\theta,x,a)$ is a known, differentiable function of the parameters $\theta$, context $x$, and action $a$; $T(y)$ denotes the sufficient statistics of the distribution; and $A(\theta,x,a)$ is the log-partition function that ensures normalization.
For exponential family models of this form, maximum likelihood estimation (MLE) can be performed efficiently by solving the associated score equation. Given a dataset $\mathcal{D}=\left\{(x_{i},a_{i},y_{i})\right\}_{i=1}^{n}$, the log-likelihood function takes the form:
$\sum_{i=1}^{n}\log h(y_{i})+\sum_{i=1}^{n}\eta(\theta,x_{i},a_{i})T(y_{i})-\sum_{i=1}^{n}A(\theta,x_{i},a_{i}).$
Taking the gradient of the log-likelihood with respect to $\theta$ and setting it to zero yields the following score equation:
$\sum_{i=1}^{n}\textbf{J}_{x_{i},a_{i}}^{T}(\theta)T(y_{i})-\sum_{i=1}^{n}\nabla_{\theta}A(\theta,x_{i},a_{i})=0.$
where $\mathbf{J}_{x_{i},a_{i}}(\theta)$ denotes the Jacobian matrix of $\eta(\theta,x_{i},a_{i})$ with respect to $\theta$. Solving this score equation yields the MLE estimate of the parameter $\theta$. For more papers about numerically computing the MLE, one could refer to some numerical computation papers  (Städler et al., 2010; Griebel et al., 2019) for details.
10 Discussion and Future Work
In this paper, we proposed a unified framework for general constrained online decision-making with uncertainty. Our framework encompasses a wide range of important applications, including constrained contextual bandits, stream-based active learning, online hypothesis testing, and online model calibration. A key theoretical contribution of this work is the generalization of the eluder dimension—from its original formulation for reward functions under squared loss to arbitrary probability divergences over general density function classes. This extension enables a principled characterization of statistical complexity in a much broader range of learning problems.
We further introduced a general algorithm that minimizes utility regret subject to stage-wise constraints. Our algorithm seamlessly integrates with any offline density estimation oracle, thereby offering a modular and widely applicable solution to constrained contextual decision-making. To support its implementation and analysis, we provided several examples illustrating how to bound the generalized eluder dimension for common density model classes, such as linear models, Gaussians, and exponential families.
Together, our contributions offer new insights into the design of reliable online algorithms under distributional uncertainty, and establish a foundation for further theoretical and algorithmic developments in this space.
Looking forward, our work opens up several promising directions. First, a natural extension is to incorporate state transitions and move toward Markov Decision Process (MDP) settings. This would enable the design of unified algorithms for general statistical tasks in MDPs, beyond the traditional reward maximization paradigm. Second, the generalized eluder dimension we introduce serves as a new lens for understanding the learnability of density model classes, and deserves further theoretical investigation. In particular, exploring how generalized eluder dimensions behave under different divergence measures for the same model class remains an open and intriguing question. Lastly, we believe it is valuable to apply our algorithm in real-world systems and empirically evaluate its practical performance across different domains.
Appendix A Useful Math Tools
Lemma A.1.
Let $\left\{X_{t}\right\}_{t=1}^{T}$ be a sequence of real-valued r.v.s. adapted to filtration $\left\{F_{t}\right\}$, $|X_{t}|\leq R$, then with probability $1-\delta$,
$\left|\sum_{t=1}^{T}X_{t}-\mathbb{E}_{t-1}[X_{t}]\right|\leq R\sqrt{8T\log(2/\delta)}.$
This lemma is called Azuma-Hoeffding inequality. For example, see [Sason, 2011].
Lemma A.2.
With the same condition in lemma A.1, for any $\eta\in(0,1/R)$,
$\sum_{t=1}^{T}X_{t}\leq\eta\sum_{t=1}^{T}\mathbb{E}_{t-1}X_{t}^{2}+\frac{\log(1/\delta)}{\eta}.$
This is called Freedman’s inequality. See proof in Agarwal et al. [2012]. Then we could have lemma A.3.
Lemma A.3.
$\left\{X_{t}\right\}$ r.v.s, adapted to filtration $F_{t}$, $0\leq X_{t}\leq R$, then with probability at least $1-\delta$,
$\sum_{t=1}^{T}X_{t}\leq\frac{3}{2}\mathbb{E}_{t-1}[X_{t}]+4R\log(2/\delta)$
$\sum_{t=1}^{T}\mathbb{E}_{t-1}[X_{t}]\leq 2\sum_{t=1}^{T}X_{t}+8R\log(2\delta^{-1}).$
Lemma A.4.
For any sequence of real-valued random variables $\left\{X_{t}\right\}_{t\leq T}$, adapted to some filtration flow $\mathcal{F}_{t}$, it holds that with probability at least $1-\delta$, for all $T^{\prime}\leq T$,
$\sum_{t=1}^{T^{\prime}}X_{t}\leq\sum_{t=1}^{T^{\prime}}\log(\mathbb{E}_{t-1}[e^{X_{t}}])+\log(1/\delta).$
The proof of lemma A.4 could be found in Foster et al. [2021]. Finally, we invoke the following result called Weyl’s inequality about the changes to eigenvalues of a Hermitian matrix that is perturbed, [Franklin, 2012].
Lemma A.5.
Assume that $A$ and $B$ are two Hermitian matrices with dimension $n$. For any matrix $M$ if we use $\lambda_{i}(M)$ to denote its $i$-th largest eigenvalue, then, $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-n}(A+B)$.
Appendix B Proofs in Section 5
Proof of lemma 5.6.
We prove this lemma by the following steps. First, if $\omega_{\mathcal{G}_{t}}(x_{t},a_{t})\geq\varepsilon$, by definition, we know that there are $g,g^{\prime}\in\mathcal{G}_{t}$ such that $\mathsf{D}(g_{x_{t},a_{t}}||g^{\prime}_{x_{t},a_{t}})>\varepsilon$. Therefore, if $(x_{t},a_{t})$ is $\varepsilon$-dependent on a subsequence $\left\{(x_{j_{1}},a_{j_{1}}),\cdots,(x_{j_{k}},a_{j_{k}}))\right\}\subset\left\{(x_{1},a_{1}),\cdots,(x_{t-1},a_{t-1})\right\}$, then
$\sum_{i=1}^{k}\mathsf{D}^{2}\left(g_{x_{j_{i}},a_{j_{i}}}||g^{\prime}_{x_{j_{i}},a_{j_{i}}}\right)\geq\varepsilon^{2}.$
Then if $(x_{t},a_{t})$ is $\varepsilon$-dependent on $H$ disjoint subsequences of $\left\{(x_{1},a_{1}),\cdots,(x_{t-1},a_{t-1})\right\}$, then
$\sum_{i=1}^{t-1}\mathsf{D}^{2}\left(g_{x_{i},a_{i}}||g^{\prime}_{x_{i},a_{i}}\right)\geq H\varepsilon^{2}.$
Since probability divergence is a metric of probability density space, by triangle inequality we have,
$\left(\sum_{i=1}^{t-1}\mathsf{D}^{2}\left(g_{x_{i},a_{i}}||g^{\prime}_{x_{i},a_{i}}\right)\right)^{1/2}\leq\left(\sum_{i=1}^{t-1}\mathsf{D}^{2}\left(g_{x_{i},a_{i}}||\widehat{g}^{t}_{x_{i},a_{i}}\right)^{1/2}\right)+\left(\sum_{i=1}^{t-1}\mathsf{D}^{2}\left(g^{\prime}_{x_{i},a_{i}}||\widehat{g}^{t}_{x_{i},a_{i}}\right)\right)^{1/2},$
$\left(\sum_{i=1}^{t-1}\mathsf{D}^{2}\left(g_{x_{i},a_{i}}||g^{\prime}_{x_{i},a_{i}}\right)\right)^{1/2}\leq 2\sqrt{r_{t}}.$
Hence, $H\leq\frac{4r_{t}}{\varepsilon^{2}}\leq\frac{4r_{T}}{\varepsilon^{2}}$. This leads to the conclusion that if $\omega_{\mathcal{G}_{t}}(x_{t},a_{t})>\varepsilon$, then $(x_{t},a_{t})$ is $\varepsilon$-dependent on fewer than $\frac{4r_{T}}{\varepsilon^{2}}$ disjoint sub-sequences of $\left\{(x_{1},a_{1}),\cdots,(x_{t-1},a_{t-1})\right\}$.
For the second step, we claim that in any context sequence $\left\{(X_{,}A_{1}),\cdots,(X_{\tau},A_{\tau})\right\}$, there is some element $(X_{j},A_{j})$ that is $\varepsilon$-dependent on at least $\tau/d-1$ disjoint subsequences of $\left\{(X_{,}A_{1}),\cdots,(X_{j-1},A_{j-1})\right\}$, where $d=\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)$. Consider any integer $H$ such that $Hd+1\leq\tau\leq(H+1)d$, we launch the following procedure to construct $H$ sequences $\left\{\text{B}_{i}\right\}_{i=1}^{H}$.
Initially, we let $\text{B}_{i}=\left\{(X_{i},A_{i})\right\}$ for all $i$. If $(X_{H+1},A_{H+1})$ is $\varepsilon$-dependent on each subsequence, then the claim is proved. Otherwise, select a subsequence such that $(X_{H+1},A_{H+1})$ is $\varepsilon$-independent and append $(X_{K+1},A_{K+1})$ to it. Repeat this process for $j>H+1$ until $(X_{j},A_{j})$ is $\varepsilon$-dependent on every subsequence or $j=\tau$. If $j=\tau$, then we have $\sum_{i=1}^{H}|B_{i}|\geq Hd$. Since each element of any subsequence $B_{i}$ is $\varepsilon$-independent of its predecessors, $|B_{i}|=d$. Thus, $(X_{\tau},A_{\tau})$ must be $\varepsilon$-dependent on each subsequence, otherwise it contradicts the assumption that $d=\text{dim}_{E}(\mathcal{F},D,\varepsilon)$.
Now, consider the $\left\{(X_{,}A_{1}),\cdots,(X_{\tau},A_{\tau})\right\}$ to be subsequence $\left\{(x_{t_{1}},a_{t_{1}})\cdots,(x_{t_{\tau}},a_{t_{\tau}})\right\}$ of $\left\{(x_{1},a_{1}),\cdots,(x_{t-1},a_{t-1})\right\}$ consisting of the elements that $\omega_{\mathcal{G}_{t}}(x_{t},a_{t})>\varepsilon$. As we have proved in the first step, $(x_{t_{j}},a_{t_{j}})$ is $\varepsilon$-dependent on fewer than $4r_{T}/\varepsilon^{2}$ subsequences of $\left\{(x_{1},a_{1}),\cdots,(x_{t_{j}-1},a_{t_{j}-1})\right\}$. Combining this with the second step that there is some $(x_{j},a_{j})$ which is $\varepsilon$-dependent on at least $\tau/d-1$ disjoint subsequences of $\left\{(x_{1},a_{1}),\cdots,(x_{j-1},a_{j-1})\right\}$, we have $\tau/d-1\leq 4r_{T}/\varepsilon^{2}$, which is
$\tau\leq(4r_{T}/\varepsilon^{2}+1)d.$
We finish the proof. ∎
Proof of lemma 5.7.
For the simplicity of notation, we write $\omega_{t}$ for $\omega_{\mathcal{G}_{t}}(x_{t},a_{t})$. We reorder the sequence such that $\omega_{i_{1}}\geq\omega_{i_{2}}\geq\cdots\geq\omega_{i_{T}}$. Then,
$\sum_{t=1}^{T}\omega_{\mathcal{G}_{t}}(x_{t},a_{t})\leq\frac{1}{T}+\sum_{t=1}^{T}\omega_{i_{t}}\mathbb{I}\left\{\omega_{i_{t}}>\alpha_{T}\right\},$
where we denote $\alpha_{T}=\frac{1}{T^{2}}$. On one hand, we know that the probability divergence is upper bounded by $C$. Since we reorder the sequence, we have,
$\omega_{i_{t}}\geq\varepsilon\Leftrightarrow\sum_{k=1}^{T}\mathbb{I}\left\{\omega_{k}>\varepsilon\right\}\geq t.$
By lemma 5.6, this can only occur when $t<(\frac{4r_{T}^{1/2}}{\varepsilon}+1)\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)$. Moreover, we have $\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)\leq\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})=d$ for $\varepsilon\geq\alpha_{T}$. Therefore, when $\omega_{i_{t}}>\varepsilon\geq\alpha_{T}$, $t\leq(\frac{4r_{T}^{1/2}}{\varepsilon}+1)$ and thus $\varepsilon\leq\sqrt{\frac{4r_{T}d}{t-d}}$. Therefore, if $\omega_{i_{t}}>\alpha_{T}$ we can conclude that $\omega_{i_{t}}\leq\min\left\{C,\sqrt{\frac{4r_{T}d}{t-d}}\right\}$. Then we have
$\sum_{t=1}^{T}\omega_{i_{t}}\mathbb{I}\left\{\omega_{i_{t}}\geq\alpha_{T}\right\}\leq Cd+\sum_{t=d+1}^{T}\sqrt{\frac{4r_{T}d}{t-d}}\leq Cd+4\sqrt{r_{T}T}.$
Finally, we combine the two terms together and use the upper bound of the probability divergence to get
$\sum_{t=1}^{T}\omega_{i_{t}}$ $\leq\min\left\{CT,\frac{1}{T}+C\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})+4\sqrt{r_{T}\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})T}\right\}$
$\leq\frac{1}{T}+\min\left\{C\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T}),CT\right\}+4\sqrt{r_{T}\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})T}.$
Appendix C Proofs in Section 7
We define $Y_{f,i}=\left({\mathcal{T}}_{1}(f(x_{i},a_{i},y))-{\mathcal{T}}_{1}(f^{*}(x_{i},a_{i},y))\right)^{2}$. We assume that the utility functional is in $[0,1]$. Then,
$Y_{f,i}^{2}\leq 4Y_{f,i}.$
Lemma C.1.
For a fixed $t\geq 2$ and fixed $\delta_{t}\in(0,1/e^{2})$, with probability at least $1-\log_{2}(t-1)\delta_{t}$, for any $f\in\mathcal{F}$,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[\left({\mathcal{T}}_{1}(f_{x_{i},a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x_{i},a_{i}})\right)^{2}|H_{i-1}\right]\leq 68\log(|\mathcal{F}|/\delta_{t})+2\sum_{i=1}^{t-1}Y_{f,i}.$
Proof of lemma C.1.
$|Y_{f,i}|\leq 1$, thus by Freedman’s Inequality A.2, with probability at least $1-\log_{2}(t-1)\delta_{t}/|\mathcal{F}|$,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[Y_{f,i}|H_{i-1}\right]-\sum_{i=1}^{t-1}Y_{f,i}\leq 4\sqrt{\sum_{i=1}^{t-1}var(Y_{f,i}|H_{i-1})\log(|\mathcal{F}|/\delta_{t})}+2\log(|\mathcal{F}|/\delta_{t}).$
Taking the union bound of $\mathcal{F}$, we get, with probability at least $1-\log_{2}(t-1)\delta_{t}$, for any $f\in\mathcal{F}$,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[Y_{f,i}|H_{i-1}\right]-\sum_{i=1}^{t-1}Y_{f,i}\leq 4\sqrt{\sum_{i=1}^{t-1}var(Y_{f,i}|H_{i-1})\log(|\mathcal{F}|/\delta_{t})}+2\log(|\mathcal{F}|/\delta_{t}).$
Since $var(Y_{f,i}|H_{i-1})\leq 4\mathbb{E}_{x_{i},a_{i}}[Y_{f,i}|H_{i-1}]$, we plug this back and complete the square method,
$\left(\sqrt{\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[Y_{f,i}|H_{i-1}\right]}-4\sqrt{\log(|\mathcal{F}|/\delta_{t})}\right)^{2}\leq 18\log(|\mathcal{F}|/\delta_{t})+2\sum_{i=1}^{t-1}Y_{f,i}.$
Therefore,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[Y_{f,i}|H_{i-1}\right]\leq 68\log(|\mathcal{F}|/\delta_{t})+2\sum_{i=1}^{t-1}Y_{f,i}.$
Proof of lemma 7.1.
For any $\delta$, define $\delta_{t}=\delta/2t^{3}$, then we have $\sum_{t=1}^{\infty}\delta_{t}\log_{2}(t-1)\leq\sum_{t}\delta/2\delta^{2}\leq\delta/2$. Therefore, after taking a union bound over $t$, with probability at least $21-\delta/$ 2,
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[Y_{f_{t},i}|H_{i-1}\right]\leq 68\log(|\mathcal{F}|t^{3}/\delta_{t})+2\sum_{i=1}^{t-1}Y_{f_{t},i}.$
Plug the definition of $Y_{f,i}$ in and we shall finish the proof. ∎
Now we can prove the following important estimation error lemma.
Lemma C.2.
Consider a measurable randomized contextual bandit algorithm that selects stochastic policy $\pi_{t}$ based on $H_{t-1}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t>K$ and every stochastic policy $\pi\in\Pi$, the expected estimation error is upper bounded by
$\left|\mathbb{E}_{x\sim Q}\left[\left\{\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi})-\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{68\log(2|\mathcal{F}|t^{3}/\delta)+4L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})}.$
Proof of lemma C.2.
First, note that ${\mathcal{T}}_{1}$ is $L_{1}$-Lipschitz continuous with respect to the TV distance, then we have for any $f^{1}$ and $f^{2}$,
$\left({\mathcal{T}}_{1}(f^{1}_{x,a})-{\mathcal{T}}_{1}(f^{2}_{x,a})\right)^{2}\leq 2L_{1}^{2}\mathsf{D}^{2}(f^{1}_{x,a}||f^{2}_{x,a}).$
Since this holds for any $x,a$, we have,
$\sum_{i=1}^{t-1}Y_{\widehat{f}^{t},i}=\sum_{i=1}^{t-1}\left({\mathcal{T}}_{1}(\widehat{f}^{t}_{x_{i},a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x_{i},a_{i}})\right)^{2}\leq 2L_{1}^{2}\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{f}^{t}_{x_{i},a_{i}}||f^{*}_{x_{i},a_{i}}).$
By 5.4, we know that with probability at least $1-\delta_{t}$, for any $f\in\mathcal{F}$,
$\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{f}^{t}_{x_{i},a_{i}}||f^{*}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta_{t}),$
Similar to lemma 7.1, for any $\delta\in(0,1)$ we set $\delta_{t}=\delta/t^{3}$, then $\sum_{t=1}^{\infty}\delta_{t}\leq\frac{\delta}{2}$, taking the union bound of $t$, we get that with probability at least $21-\delta/$ 2,
$\sum_{i=1}^{t-1}\mathsf{D}^{2}(\widehat{f}^{t}_{x_{i},a_{i}}||f^{*}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3}),\ \forall t,f^{*}\in\mathcal{F}.$
By some calculation, we have
$\mathbb{E}_{x_{i}\sim Q,a_{i}\sim\pi_{i}}\left[\left({\mathcal{T}}_{1}(f_{x_{i},a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x_{i},a_{i}})\right)^{2}|H_{i-1}\right]$
$=$ $\mathbb{E}_{x\sim Q,a\sim\pi_{i}(\cdot|x)}\left[\left({\mathcal{T}}_{1}(f_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\big{|}H_{i-1}\right]$
$=$ $\mathbb{E}_{x\sim Q}\left[\sum_{a}\pi_{i}(a|x)\left({\mathcal{T}}_{1}(f_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}|H_{i-1}\right]$
$=$ $\mathbb{E}_{x\sim Q}\left[\sum_{a}\pi_{i}(a|x)\left({\mathcal{T}}_{1}(f_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\right]$
$=$ $\mathbb{E}_{x\sim Q,a\sim\pi_{i}}\left[\left({\mathcal{T}}_{1}(f_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\right]$
The first equality is due to the fact that $a_{i}\sim\pi_{i}$ and $\pi_{i}$ is completely determined by $H_{i-1}$; the second is due to the independence of $x_{i}$ and $H_{i-1}$; the third one is because $\left(\tilde{{\mathcal{T}}_{1}}(f_{x,\pi_{i}})-\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi_{i}})\right)^{2}$ is related to $H_{i-1}$ only through $\pi_{i}$.
We notice that for any policy $\pi$ and any $f$,
$\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}\left({\mathcal{T}}_{1}(f_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\right]$
$=$ $\sum_{i=1}^{t-1}\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\mathbb{I}\left\{a=a_{i}\right\}\left({\mathcal{T}}_{1}(f_{x,a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x,a_{i}})\right)^{2}\right]$
Then we apply lemma 7.1 to get
$\sum_{i=1}^{t-1}\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\mathbb{I}\left\{a=a_{i}\right\}\left({\mathcal{T}}_{1}(f^{t}_{x,a_{i}})-{\mathcal{T}}_{1}(f^{*}_{x,a_{i}})\right)^{2}\right]\leq 68\log(|\mathcal{F}|t^{3}/\delta)+2\sum_{i=1}^{t-1}Y_{f_{t},i}.$
Then for $t>K$ and any sequence $\left\{f_{t}\right\}_{t>K}$, by the Cauchy-Schwarz Inequality, we have that with probability at least $21-\delta/$ 2,
$=$ $\left|\mathbb{E}_{x\sim Q,a\sim\pi}\left[{\mathcal{T}}_{1}(f^{t}_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right]\right|$
$=$ $\left|\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[{\mathcal{T}}_{1}(f^{t}_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right]\right|$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}\left({\mathcal{T}}_{1}(f^{t}_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a}))\right)^{2}\right]}$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{68\log(|\mathcal{F}|t^{3}/\delta)+2\sum_{i=1}^{t-1}Y_{f^{t},i}}$
The first inequality is by Cauchy-Schwarz Inequality and the second is by lemma 7.1.
Taking $f^{t}=\widehat{f}^{t}$ in the above inequality and using our offline density estimation guarantee, we have that with probability at least $1-\delta$,
$\left|\mathbb{E}_{x\sim Q}\left[\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi})-\widetilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi})\right]\right|$
$=$ $\left|\mathbb{E}_{x\sim Q,a\sim\pi}\left[{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right]\right|$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{68\log(2|\mathcal{F}|t^{3}/\delta)+4L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})}.$
Lemma C.3.
Consider a measurable randomized contextual bandit algorithm that selects a stochastic policy $\pi_{t}$ based on $H_{t-1}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t>K$ and every stochastic policy $\pi\in\Pi$, the expected estimation error is upper bounded by
$\left|\mathbb{E}_{x\sim Q}\left[\left\{\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi})-\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{K\beta_{t}}{t},$
where
$\beta_{t}=\sqrt{\frac{(34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3}))t}{K}}.$
Proof of lemma C.3.
We first apply lemma C.2 and use the AM-GM inequality to get:
$\left|\mathbb{E}_{x\sim Q}\left[\left\{{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,\pi})-{\mathcal{T}}_{1}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\frac{1}{2}\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})}{\beta_{t}}.$
Then we focus on the term $\frac{34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})}{\beta_{t}}$. Recall that we set
$\beta_{t}=\sqrt{\frac{(34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3}))t}{K}},$
Then, we know that
$\frac{34\log(2|\mathcal{F}|t^{3}/\delta)+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\delta/2t^{3})}{\beta_{t}}=\frac{K\beta_{t}}{t}.$
By plugging this back and noticing that we enlarge the coefficient of the first term from $1/2$ to $1$, we finish the proof. ∎
To bound the first term in lemma C.3, we need the following potential lemma.
Lemma C.4.
Let $\left\{a_{t}\right\}_{t=1}^{T}$ be the action sequence that at first $K$ rounds, we explore every arm regardless of the context. After the first $t$ rounds, we place no restrictions on the sequence. Then for any $T>K$, we have,
$\sum_{t=K+1}^{T}\left[\frac{1}{\sum_{j=1}^{t-1}\mathbb{I}\left\{a_{t}=a_{j}\right\}}\right]\leq K+K\log(\frac{T}{K})$
Proof of lemma C.4.
We prove this by direct algebra,
$\sum_{t=K+1}^{T}\frac{1}{\sum_{j=1}^{t-1}\mathbb{I}\left\{a_{t}=a_{j}\right\}}\leq\sum_{a\in\mathcal{A}}\sum_{i=1}^{\sum_{t=1}^{T}\mathbb{I}\left\{a_{t}=a\right\}}\frac{1}{i}\leq\sum_{a\in\mathcal{A}}(1+\log(\sum_{t=1}^{T}\mathbb{I}\left\{a_{t}=a\right\}))\leq K+K\log(T/K).$
Proof of lemma 7.3.
We use the statistical guarantee of our offline density estimation oracle. By assumption, with probability $21-\delta/$ 2,
$\sum_{i=1}^{t-1}\mathsf{D}^{2}(g^{*}_{x_{i},a_{i}}||\widehat{g}^{t}_{x_{i},a_{i}})\leq\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{G},t,\frac{\delta}{2t^{3}}),$
uniformly for all $t$. Because we already know about the property of the safe action $a_{0}$, we have that ${\mathcal{T}}_{2}(g^{*}_{x,a_{0}})=c_{0}$. Therefore, we know that with probability at least $1-\delta$, $g^{*}\in\tilde{C}_{\mathcal{G}}(t,\delta)$ holds uniformly for all $t$. ∎
Proof of lemma 7.5.
First of all, we have with probability at least $1-\delta$,
$\tilde{V}_{\delta}^{t}(x,\pi)\leq\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,\pi})+\max_{g,g^{\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})-\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})\right\}.$
This is because $g^{*}\in\tilde{C}_{\mathcal{G}}(t,\delta)$ with high probability. Then,
$\tilde{V}_{\delta}^{t}(x,\pi)=$ $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,\pi})+\max_{g\in\tilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})-{\mathcal{T}}_{2}(g^{*}_{x,\pi})\right\}$
$\leq$ $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,\pi})+\max_{g,g^{\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})-\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})\right\}$
From lemma C.2, we also know that with high probability, for any policy $\pi$
$\mathbb{E}_{x\sim Q,a\sim\pi}{\mathcal{T}}_{1}(f^{*}_{x,a})\leq\mathbb{E}_{x\sim Q,a\sim\pi}{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,a})+\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{K\beta_{t}}{t}.$
According to Algorithm 1, our algorithm selects
[…]
$+\alpha_{r}\mathbb{E}_{x\sim Q,a\sim\pi}\max_{g^{\prime},g^{\prime\prime}\in\widetilde{C}_{\mathcal{G}}(t,\delta)}{\mathcal{T}}_{2}(g^{\prime}_{x_{t},a})-{\mathcal{T}}_{2}(g^{\prime\prime}_{x_{t},a})\Bigg{\}}.$
Then we prove that with high probability, our adjusted policy upper confidence bound has,
$\mathbb{E}_{x\sim Q}\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi_{t}})+\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{K\beta_{t}}{t}$
$+\alpha_{r}\mathbb{E}_{x\sim Q,a\sim\pi}\max_{g^{\prime},g^{\prime\prime}\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\left\{{\mathcal{T}}_{2}(g^{\prime}_{x,a})-{\mathcal{T}}_{2}(g^{\prime\prime}_{x,a})\right\}\geq\mathbb{E}_{x}\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,\pi^{*}}),$
$\pi^{*}$ is the optimal feasible policy, and $\pi_{t}$ is the optimal upper confidence bound policy that we choose in round $t$.
If $\pi^{*}(x)\in\tilde{\Pi}^{t}_{\delta}(x)$, then we could directly apply lemma C.2 to get the result. Now we focus on the case where $\pi^{*}(x)\notin\tilde{\Pi}^{t}_{\delta}(x)$. Let $\pi_{0}(x)=a_{0}$ be the policy that always applies safe action. By definition, ${\mathcal{T}}_{2}(g^{*}_{x,a_{0}})=c_{0}$. Consider a mixed policy, given any $x$, $\tilde{\pi}_{t}(\cdot|x)=\gamma_{t}^{x}\pi^{*}_{t}(\cdot|x)+(1-\gamma_{t}^{x})\pi_{0}(\cdot|x)$. Thus, we define $\gamma_{t}^{x}$ to be the smallest number such that under this mixed policy, $\max_{g\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\tilde{\pi}_{t}})=\tau$. Recall that $\widetilde{V}^{t}_{\delta}(x,\pi)=\max_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})$, for the optimal stochastic policy $\pi^{*}$, it is obvious that $\pi^{*}(\cdot|x)$ is also optimal for every context $x$, let $\widetilde{g}^{x,t}=\mathop{\mathrm{argmax}}_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi^{*}})$.
Because when we apply the safe policy $\pi_{0}=a_{0}$, we will always encounter a cost $c_{0}$ regardless of the context, then
$\widetilde{g}^{x,t}=\mathop{\mathrm{argmax}}_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\gamma\pi^{*}+(1-\gamma)\pi_{0}}).$
Also,
$\widetilde{V}^{t}_{\delta}(x,\gamma\pi^{*}+(1-\gamma)\pi_{0})=\gamma\tilde{{\mathcal{T}}_{2}}(\widetilde{g}^{x,t}_{x,\pi^{*}})+(1-\gamma)c_{0}.$
This indicates that there is a number $\gamma_{t}^{x}$ such that
$\widetilde{V}^{t}_{\delta}(x,\gamma_{t}^{x}\pi^{*}+(1-\gamma_{t}^{x})\pi_{0})=\tau.$
By definition of $\widetilde{V}^{t}_{\delta}$, we have
$\gamma_{t}^{x}\widetilde{V}^{t}_{\delta}(x,\pi^{*})+(1-\gamma_{t}^{x})c_{0}=\tau.$
$\gamma_{t}^{x}=\frac{\tau-c_{0}}{\widetilde{V}^{t}_{\delta}(x,\pi^{*})-c_{0}}\geq\frac{\tau-c_{0}}{\tau-c_{0}+\max_{g,g^{\prime}\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\left\{\tilde{{\mathcal{T}}_{2}}(g_{x,\pi^{*}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi^{*}})\right\}}.$
Since $\pi_{t}(\cdot|x)$ and $\widetilde{\pi}_{t}(\cdot|x)$ are both feasible at $x$, we have that
$\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi_{t}})+\beta_{t}\mathbb{E}_{a\sim\pi_{t}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi_{t}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi_{t}})$
$\geq$ $\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\widetilde{\pi}_{t}})+\beta_{t}\mathbb{E}_{a\sim\widetilde{\pi}_{t}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\widetilde{\pi}_{t}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\widetilde{\pi}_{t}})$
$=$ $\gamma_{t}^{x}\left(\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}})+\beta_{t}\mathbb{E}_{a\sim\pi^{*}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi^{*}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi^{*}})\right)$
$+$ $(1-\gamma_{x}^{t})\left(r_{0}+\beta_{t}\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a_{0}=\widetilde{a}_{t,j}\right\}+1}\right)$
$\geq$ $\gamma_{x}^{t}\left(\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}})+\beta_{t}\mathbb{E}_{a\sim\pi^{*}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}C_{1}\right)+(1-\gamma_{x}^{t})r_{0},$
where $C_{1}=\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi^{*}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi^{*}})$. Recall that we could set $\gamma_{t}^{x}=\frac{\tau-c_{0}}{\tau-c_{0}+C_{1}}$, we want to set our $\alpha_{r}$ to have,
$\gamma_{x}^{t}\left(\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}})+\beta_{t}\mathbb{E}_{a\sim\pi^{*}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}C_{1}\right)+(1-\gamma_{x}^{t})r_{0}\geq\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}}).$
By some algebra, we could set $\alpha_{r}=\frac{1-r_{0}}{\tau-c_{0}}$ and it could be satisfied.
Finally, we conclude that, for any $x$, we select $\pi_{t}$ according to our algorithm, then,
$\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi_{t}})+\beta_{t}\mathbb{E}_{a\sim\pi_{t}}\left[\frac{1}{\sum_{j=K+1}^{t-1}\mathbb{I}\left\{a=\widetilde{a}_{t,j}\right\}+1}\right]+\alpha_{r}\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi_{t}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi_{t}})\geq\tilde{{\mathcal{T}}_{1}}(\widehat{f}^{t}_{x,\pi^{*}}),$
where $\pi^{*}$ is the optimal feasible stochastic policy given $x$ and $\pi_{t}$ is the stochastic policy at round $t$. So we finish the proof. ∎
With all these lemmas equipped, now we can prove our main theoretical guarantee.
Proof of theorem 7.6.
First, by lemma C.2, we have
$\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi^{*}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right]-\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right]$
$\leq$ $\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi^{*}(\cdot|x_{t})}{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,a})+\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{K\beta_{t}}{t}-\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right].$
Then we apply lemma 7.5 to get,
$\sum_{t=1}^{T}\left\{\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi^{*}(\cdot|x_{t})}{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,a})+\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]-\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right]\right\}$
$\leq$ $\sum_{t=1}^{T}\left\{\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x)}{\mathcal{T}}_{1}(\widehat{f}_{t}(x,a_{t},y))-\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right]\right\}+\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]$
$+\alpha_{r}\sum_{t=1}^{T}\mathbb{E}_{x\sim Q}\left[\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}{\mathcal{T}}_{2}(g^{\prime}_{x,\pi_{t}})-{\mathcal{T}}_{2}(g^{\prime\prime}_{x,\pi_{t}})\right]$
For the first term, again we apply lemma C.2,
$\sum_{t=1}^{T}\left\{\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x)}{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,\pi_{t}})-\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}(\cdot|x_{t})}\left[{\mathcal{T}}_{1}(f^{*}_{x_{t},a_{t}})\right]\right\}$
$\leq$ $\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\sum_{t=1}^{T}\frac{K\beta_{t}}{t}.$
Combining all these parts together, we can upper bound the regret by
$2\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+2\sum_{t=1}^{T}\frac{K\beta_{t}}{t}+\alpha_{r}\sum_{t=1}^{T}\mathbb{E}_{x\sim Q}\left[\max_{g^{\prime},g^{\prime\prime}\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g^{\prime}_{x,\pi_{t}})-\tilde{{\mathcal{T}}_{2}}(g^{\prime\prime}_{x,\pi_{t}})\right].$
The first two terms are already well bounded. To bound the third term, we need to use the properties of generalized eluder dimension.
Specifically, we apply lemma 5.7 to get
$\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}}\left[\max_{g^{\prime},g^{\prime\prime}\in\widetilde{C}_{\mathcal{G}}(t,\delta)}{\mathcal{T}}_{2}(g^{\prime}_{x,a_{t}})-{\mathcal{T}}_{2}(g^{\prime\prime}_{x,a_{t}})\right]$
$\leq$ $\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t}}\left[C\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},\frac{1}{T^{2}}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}\right]+1$
$=$ $C\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},\frac{1}{T^{2}}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}+1.$
Plug this bound in, we finish the proof. ∎
C.1 Proofs in Section 7.2
Proof of corollary 7.8.
Since we have two density model classes. We focus on the utility density class $\mathcal{F}$. The discussion of the constraint model class $\mathcal{G}$ is similar.
From the result on the covering of $d$-dimensional balls, the covering number of a $d$-dimensional ball with radius $\frac{\Delta}{2}$ and discretization error $\frac{1}{L_{2}t}$ is bounded by $(1+\Delta L_{2}t)^{d}$. So there exists a set $V_{t}$ of size no more than $(1+\Delta L_{2}t)^{d}+1\leq(2+\Delta L_{2}t)^{d}$ that contains the true $\theta^{*}$ and satisfies
$\forall\theta\in\Theta,\exists v\in V_{t},\text{s.t.}||\theta-v||_{2}\leq\frac{1}{L_{2}t}.$
$\log|V_{t}|\leq d\log(2+\Delta L_{2}t)$. For any $f_{\theta}\in\mathcal{F}$, for any $x\in\mathcal{X}$, $a\in\mathcal{A}$, set $v$ to be the closest point to $\theta$ in $V_{t}$, we have
$\left({\mathcal{T}}_{1}(f_{\theta,x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\leq 2\left({\mathcal{T}}_{1}(f_{\theta,x,a})-{\mathcal{T}}_{1}(f_{v,x,a})\right)^{2}+2\left({\mathcal{T}}_{1}(f_{v,x,a})-{\mathcal{T}}_{1}(f_{\theta^{*},x,a})\right)^{2}.$
We first use the Lipschitz continuity of the functional ${\mathcal{T}}_{1}$ and then use the parametric assumption of the density class to get
$2\left({\mathcal{T}}_{1}(f_{\theta,x,a})-{\mathcal{T}}_{1}(f_{v,x,a})\right)^{2}\leq 2L_{\mathsf{D}}^{2}L_{2}^{2}||\theta-v||_{2}^{2}.$
Then we could prove the following result similar to lemma C.1. First, we view $V_{t}$ as a finite density class and apply lemma C.1 to get with probability at least $1-\log_{2}(t-1)\delta_{t}$:
$2\mathbb{E}_{x_{i},a_{i}}\left[\left({\mathcal{T}}_{1}(f_{v,x,a})-{\mathcal{T}}_{1}(f_{\theta^{*},x,a})\right)^{2}|H_{i-1}\right]\leq 136\log(|V_{t}|/\delta_{t})+4\sum_{i=1}^{t-1}Y_{f,i}.$
Combining all these together, we have with probability $1-\log_{2}(t-1)\delta_{t}$,
$\mathbb{E}_{x_{i},a_{i}}\left[\left({\mathcal{T}}_{1}(f_{\theta,x,a})-{\mathcal{T}}_{1}(f^{*}_{x,a})\right)^{2}\right]\leq 136\log(|V_{t}|/\delta_{t})+4\sum_{i=1}^{t-1}Y_{f,i}+2L_{\mathsf{D}}^{2},$
uniformly over all $f_{\theta}$.
Then, similar to lemma 7.1, we set $\delta_{t}=\frac{\delta}{t^{3}}$ and apply the union bound over $t$ to get with probability at least $21-\delta/$ 2
$\sum_{i=1}^{t-1}\mathbb{E}_{x_{i},a_{i}}\left[({\mathcal{T}}_{1}(f^{t}_{x_{i},a_{i}}-{\mathcal{T}}_{1}(f_{x_{i},a_{i}})))^{2}|H_{i-1}\right]\leq 136\left(d\log(2+\Delta L_{2}t)+\log\frac{2t^{3}}{\delta}\right)+2L_{\mathsf{D}}^{2}+4\sum_{i=1}^{t-1}Y_{f_{t},i},$
uniformly over all $t\geq 2$ and sequence $f^{2},f^{3},\cdots\in\mathcal{F}$. Then similar to the proof of lemma C.2, we could show that
$\left|\mathbb{E}_{x\sim Q}\left[\left\{{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,\pi}-{\mathcal{T}}_{1}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\sqrt{\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{1}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]}\sqrt{136\left(d\log(2+\Delta L_{2}t)+\log(\frac{2t^{3}}{\delta})\right)+2L_{\mathsf{D}}^{2}+4L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\frac{\delta}{2t^{3}})}.$
Setting
$\beta_{t}=\sqrt{\frac{\left\{72\left(d\log(2+\Delta L_{2}t)+\log(\frac{2t^{3}}{\delta})\right)+L_{\mathsf{D}}^{2}+2L_{1}^{2}\texttt{Est}_{\mathsf{D}}^{\mathsf{Off}}(\mathcal{F},t,\frac{\delta}{2t^{3}})\right\}t}{K}},$
we have,
$\left|\mathbb{E}_{x\sim Q}\left[\left\{{\mathcal{T}}_{1}(\widehat{f}^{t}_{x,\pi})-{\mathcal{T}}_{1}(f^{*}_{x,\pi})\right\}\right]\right|$
$\leq$ $\mathbb{E}_{x\sim Q,a\sim\pi,a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+\frac{K\beta_{t}}{t}.$
Then we follow the proof of theorem 7.6 to get:
$\text{Reg}(T)\leq$ $2\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim Q,a_{t}\sim\pi_{t},a_{i}\sim\pi_{i}}\left[\frac{\beta_{t}}{\sum_{i=1}^{t-1}\mathbb{I}\left\{a=a_{i}\right\}}\right]+2\sum_{t=1}^{T}\frac{K\beta_{t}}{t}$
$+\alpha_{r}\left[2\min\left\{\text{dim}_{E}(\mathcal{G},\mathsf{D},\frac{1}{T^{2}}),T\right\}+4\sqrt{\text{dim}_{E}(\mathcal{G},\mathsf{D},1/T^{2})T}+1\right].$
Plugging in the value of $\beta_{t}$ and notice that $\beta_{t}\leq\beta_{T}$, we finish the proof.
Appendix D Proofs in Section 8
Proof of theorem 8.1.
First, because we consider $L^{2}$ divergence, then
$D(f_{x,a}||f^{\prime}_{x,a})=(\theta-\theta^{\prime})^{T}\int_{\Omega}\phi_{x,a}(y)\phi^{T}_{x,a}(y)dy(\theta-\theta^{\prime})$
if $w_{k}\geq\varepsilon^{\prime}$, then by definition,
$w_{k}\leq\max\left\{\sqrt{(\theta-\theta^{\prime})^{T}m_{k}(\theta-\theta^{\prime})}:(\theta-\theta^{\prime})^{T}M_{k}(\theta-\theta^{\prime})\leq(\varepsilon^{\prime})^{2},(\theta-\theta^{\prime})^{T}I_{d}(\theta-\theta^{\prime})\leq(2R)^{2}\right\}$
Denote $\rho=\theta-\theta^{\prime}$, the right-hand side is then smaller than
$\max\left\{\sqrt{(\theta-\theta^{\prime})^{T}m_{k}(\theta-\theta^{\prime})}:(\theta-\theta^{\prime})^{T}(M_{k}+\lambda I_{d})(\theta-\theta^{\prime})\leq(\varepsilon^{\prime})^{2}+\lambda(2R)^{2}\right\}.$
The solution of this optimization problem is $\sqrt{\lambda_{\max}\left((M_{k}+\lambda I_{d})^{-1}m_{k}\right)((\varepsilon^{\prime})^{2}+\lambda(2R)^{2})}$, thus we get,
$\lambda_{\max}\left((M_{k}+\lambda I_{d})^{-1}m_{k}\right)\geq\frac{(\varepsilon^{\prime})^{2}}{(\varepsilon^{\prime})^{2}+4R^{2}\lambda}.$
Then by Weyl’s inequality lemma A.5, we have,
$\lambda_{\max}(M_{k}+\lambda I_{d})$ $=\lambda_{\max}\left((M_{k-1}+\lambda I_{d})(I_{d}+(M_{k-1}+\lambda I_{d})^{-1}m_{k-1})\right)$
$\geq\lambda_{\min}(M_{k-1}+\lambda I_{d})(1+\frac{(\varepsilon^{\prime})^{2}}{(\varepsilon^{\prime})^{2}+4R^{2}\lambda})$
$\geq$ $(1+\frac{(\varepsilon^{\prime})^{2}}{(\varepsilon^{\prime})^{2}+4R^{2}\lambda})\alpha\lambda_{\max}(M_{k-1}+\lambda I_{d})$
We set $\lambda=\frac{(2-1/\alpha)(\varepsilon^{\prime})^{2}}{8R^{2}(1/\alpha-1)}$ to get,
$(1+\frac{(\varepsilon^{\prime})^{2}}{(\varepsilon^{\prime})^{2}+4R^{2}\lambda})\alpha=3\alpha-2\alpha^{2}>1,$
Thus, we iterate this formula to get,
$\lambda_{\max}(M_{k}+\lambda I_{d})\geq(3\alpha-2\alpha^{2})^{k-1}\lambda,$
On the other hand,
$\lambda_{\max}(M_{k}+\lambda I_{d})\leq\lambda+kS^{2}.$
Then we must have,
$\lambda+(k-1)S^{2}\geq(3\alpha-2\alpha^{2})^{k-1}\lambda\Leftrightarrow(3\alpha-2\alpha^{2})^{k-1}\leq 1+\frac{S^{2}}{\lambda}(k-1).$
Manipulating this inequality, we define the following quantity: $B(x,\beta)=\max\left\{B:x^{B}\leq\beta B+1\right\}$, where $x>1$. $B\log x\leq\log(1+\beta)+\log B$. Since $\log(x)\geq\frac{x-1}{x}$ for $x>1$, using the transformation of $z=B[\frac{x-1}{x}]$, we have
$z\leq\log(1+\beta)+\log z+\log\frac{x}{x-1}\leq\log(1+\beta)+z/e+\log\frac{x}{x-1}.$
Thus, $z\leq\frac{e}{e-1}\left(\log(1+\beta)+\log\frac{x}{x-1}\right)$.
Therefore, $B\leq\frac{x}{x-1}\frac{e}{e-1}\left(\log(1+\beta)+\log\frac{x}{x-1}\right)$. Plug $x=3\alpha-2\alpha^{2}$ and $\beta=\frac{S^{2}}{\lambda}$ in, we shall get the following result.
$\text{dim}_{E}(\mathcal{F},\mathsf{D}_{L^{2}},\varepsilon)\leq\frac{3\alpha-2\alpha^{2}}{3\alpha-2\alpha^{2}-1}\frac{e}{e-1}(\log(1+\frac{8R^{2}S^{2}(1/\alpha-1)}{(2-1/\alpha)\varepsilon^{2}})+\log(\frac{3\alpha-2\alpha^{2}}{3\alpha-2\alpha^{2}-1})).$
So we finish the proof.
Proof of theorem 8.2.
If $w_{k}\geq\varepsilon^{\prime}$ then
$\varepsilon^{\prime}\leq w_{k}\leq\overline{\Delta}\max\left\{(\theta-\theta^{\prime})^{T}\phi_{x_{k},a_{k}}:(\theta-\theta^{\prime})^{T}\sum_{i=1}^{k-1}\phi_{x_{i},a_{i}}\phi^{T}_{x_{i},a_{i}}(\theta-\theta^{\prime})\leq(\varepsilon^{\prime})^{2}||\theta-\theta^{\prime}||_{2}\leq 2\right\}.$
Then we define $V_{k}=\sum_{i=1}^{k-1}+\phi_{x_{i},a_{i}}\phi^{T}_{x_{i},a_{i}}+\lambda I$, $\lambda=(\frac{\varepsilon^{\prime}}{2})^{2}$. we solve the optimization problem to get
$\overline{\Delta}\sqrt{2(\varepsilon^{\prime})^{2}}||\phi_{k}||_{V_{k}^{-1}}\geq\varepsilon^{\prime}.$
Thus $\phi_{k}^{T}V_{k}^{-1}\phi_{k}\geq\frac{1}{2\overline{\Delta}}$.
For the second step, if we have a sequence $\left\{(x_{i},a_{i})\right\}_{i=1}^{k}$ such that $w_{i}\geq\varepsilon^{\prime}$ for each $i<k$, i.e., $k\leq\text{dim}_{E}(\mathcal{F},D_{H},\varepsilon)$ , by matrix determinant lemma, we have
$\det(V_{k})=\det(V_{k-1})(1+\phi^{T}_{k}V_{k}^{-1}\phi_{k})\geq(1+\frac{1}{2\overline{\Delta}})\det(V_{k-1})\geq\cdots\geq\lambda^{d}\left(1+\frac{1}{\overline{\Delta}}\right)^{k-1}.$
On the other hand, we know that $\det V_{k}\leq(\frac{trace(V_{k})}{d})^{d}\leq\left(d+\frac{k-1}{d}\right)^{d}.$
Therefore, we must have,
$\left(1+\frac{1}{2\overline{\Delta}}\right)^{\frac{k-1}{d}}\leq\frac{1}{\lambda}[\frac{k-1}{d}]+1.$
Following the previous subsection, let $B(x,\alpha)=\max\left\{B:(x+1)^{B}\leq\alpha B+1\right\}$, from $(x+1)^{B}\leq\alpha B+1$, we get
$\log(1+x)B\leq\log(1+\alpha)+\log B.$
By applying change of variable $y=B[\frac{1+x}{x}]$, we have
$y\leq\log(1+\alpha)+\log(1+x/x)+\log y\leq\log(1+\alpha)+\log\frac{1+x}{x}+y/e$
Thus,
$y\leq\frac{e}{e-1}(\log(1+\alpha)+\log(\frac{1+x}{x})),$
so $B(x,\alpha)\leq\frac{1+x}{x}\left(\frac{e}{e-1}(\log(1+\alpha)+\log(\frac{1+x}{x}))\right)$. We set $x=\frac{1}{2\overline{\Delta}}$ and $\alpha=\frac{1}{\lambda}=\frac{4}{(\varepsilon^{\prime})^{2}}$ and get
$(k-1)/d\leq(1+2\overline{\Delta})\frac{e}{e-1}\left(\log(1+\frac{4}{(\varepsilon^{\prime})^{2}})+\log(1+2\overline{\Delta})\right).$
So
$k-1\leq d\left((1+2\overline{\Delta})\frac{e}{e-1}\left(\log(1+\frac{4}{(\varepsilon^{\prime})^{2}})+\log(1+2\overline{\Delta})\right)\right)$
Proof of theorem 8.3.
We focus on computing the Hellinger distance. By direct calculation, we have,
$D_{H}^{2}(f^{1}_{x,a}||f^{2}_{x,a})=1-\exp\left\{A(\frac{\eta_{1}+\eta_{2}}{2})-\frac{1}{2}(A(\eta_{1})+A(\eta_{2}))\right\}.$
We apply Taylor expansion of $A(\eta)$ at point $\eta_{m}=\frac{\eta+\eta^{\prime}}{2}$.
$A(\eta)=A(\frac{\eta_{1}+\eta_{2}}{2})+(\eta-\eta_{m})^{T}\nabla A(\eta_{m})+\frac{1}{2}(\eta-\eta_{m})^{T}\nabla^{2}A(\xi)(\eta-\eta_{m}),$
for some $\xi$. We plug $\eta=\eta_{1}$ and $\eta=\eta_{2}$ in respectively to get
$-\frac{1}{8}\overline{\lambda}||\eta_{1}-\eta_{2}||_{2}^{2}\leq A((\eta_{1}+\eta_{2})/2)-\frac{1}{2}A(\eta_{1})-\frac{1}{2}A(\eta_{2})\leq-\frac{1}{8}\underline{\lambda}||\eta_{1}-\eta_{2}||_{2}^{2}.$
Therefore,
$1-\exp\left\{-\frac{1}{8}\underline{\lambda}||\eta_{1}-\eta_{2}||_{2}^{2}\right\}\leq D_{H}^{2}(f^{1}_{x,a}||f^{2}_{x,a})\leq 1-\exp\left\{-\frac{1}{8}\overline{\lambda}||\eta_{1}-\eta_{2}||_{2}^{2}\right\}.$
Again, because we know that $W$ is in a compact set and $\phi$ has bounded norm, then we know that there is some constant $c$ such that $cx\leq 1-e^{-x}\leq x$.
Now, if we have $w_{k}\geq\varepsilon^{\prime}$, i.e.,
$\frac{\overline{\lambda}}{8}\max_{\rho=(W_{1}-W_{2})\phi_{k},||W_{1}||\leq\beta,||W_{2}||\leq\beta}\left\{\rho^{T}\rho:\frac{c\underline{\lambda}}{8}\sum_{i=1}^{k-1}\phi_{i}^{T}(W_{1}-W_{2})^{T}(W_{1}-W_{2})\phi_{i}\leq(\varepsilon^{\prime})^{2}\right\}$
$\geq$ $w_{k}^{2}=D_{H}^{2}(f^{1}_{x_{k},a_{k}}||f^{2}_{x_{k},a_{k}})\geq(\varepsilon^{\prime})^{2},$
where we write $\phi_{i}=\phi_{x_{i},a_{i}}$.
Furthermore, for the optimization problem
$\frac{\overline{\lambda}}{8}\max_{\rho=(W_{1}-W_{2})^{T}\phi_{k},||W_{1}||\leq\beta||W_{2}||\leq\beta}\left\{\rho^{T}\rho:\frac{c\underline{\lambda}}{8}\sum_{i=1}^{k-1}\phi_{i}^{T}(W_{1}-W_{2})^{T}(W_{1}-W_{2})\phi_{i}\leq(\varepsilon^{\prime})^{2}\right\},$
by denoting $W_{1}-W_{2}$ as $W$ and $X=W^{T}W$, we have $\sum_{i=1}^{k-1}\phi_{i}^{T}(W_{1}-W_{2})^{T}(W_{1}-W_{2})\phi_{i}=tr(X\sum_{i=1}^{k-1}\phi_{i}\phi_{i}^{T})$. Similarly, we have $\phi_{k}^{T}(W_{1}-W_{2})^{T}(W_{1}-W_{2})\phi_{k}=tr(\phi_{k}\phi_{k}^{T}X)$. Therefore,
$\max_{\rho=(W_{1}-W_{2})^{T}\phi_{k},||W_{1}||\leq\beta||W_{2}||\leq\beta}\left\{\rho^{T}\rho:\frac{c\underline{\lambda}}{8}\sum_{i=1}^{k-1}\phi_{i}^{T}(W_{1}-W_{2})^{T}(W_{1}-W_{2})\phi_{i}\leq(\varepsilon^{\prime})^{2}\right\}$
$\leq$ $\max_{X\succeq 0}\left\{tr(\phi_{k}\phi_{k}^{T}X):tr\left(X\left(\sum_{i=1}^{k-1}\phi_{i}\phi_{i}^{T}+uI\right)\right)\leq\left(\frac{8}{c\underline{\lambda}}+\kappa\right)(\varepsilon^{\prime})^{2}\right\}.$
where $\kappa(\varepsilon^{\prime})^{2}=2u\beta$.
By solving this optimization problem, we get that the optimal value of the RHS optimization problem is $\left(\frac{8}{c\underline{\lambda}}+\kappa\right)(\varepsilon^{\prime})^{2}\phi_{k}^{T}\left(\sum_{i=1}^{k-1}\phi_{i}\phi_{i}^{T}+uI\right)^{-1}\phi_{k}$. Thus, denoting $V_{k}=\sum_{i=1}^{k-1}\phi_{i}\phi_{i}^{T}+uI$ and combining with $w_{k}\geq\varepsilon^{\prime}$, we have,
$||\phi_{k}||_{V_{k}^{-1}}\geq\frac{1}{\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}}.$
Now if we have a sequence $\left\{(x_{i},a_{i})\right\}_{i=1}^{k}$ such that $w_{i}\geq\varepsilon^{\prime}$ for all $i<k$, then by matrix determinant lemma
$\det(V_{k})\geq u^{d}(1+\frac{1}{\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}})^{k-1}.$
On the other hand, we have
$\det(V_{k})\leq\left(\frac{k-1}{d}+u\right)^{d}.$
Thus we must have
$u^{d}(1+\frac{1}{\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}})^{k-1}\leq\left(\frac{k-1}{d}+u\right)^{d}.$
Similar to previous section, this implies
$\left(1+\frac{1}{\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}}\right)^{\frac{k-1}{d}}\leq\frac{1}{u}[\frac{k-1}{d}]+1.$
This indicates that
$k-1\leq d\left(\left(1+\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}\right)\frac{e}{e-1}\left(\log(1+\frac{2\beta}{\kappa(\varepsilon^{\prime})^{2}})\right)+\log\left(1+\sqrt{\kappa+\frac{8}{c\underline{\lambda}}}\right)\right).$
We set $u=\frac{(\varepsilon^{\prime})^{2}}{4\beta}$ and $\kappa=\frac{1}{2}$ and we finish the proof. ∎
Appendix E Proofs in Section 9
Proofs in theorem 9.1.
Recall that in theorem 8.2, we show that the Hellinger distance between $f_{\theta,x,a}(x,a,y)$ and $f_{\theta^{\prime},x,a}$ is bounded by $\mathcal{O}(|(\theta-\theta^{\prime})\phi(x,a)|^{2})$. By some algebra, we have that
$\sum_{i=1}^{n}\left((\theta^{*})^{T}\phi_{x_{i},a_{i}}-(\widehat{\theta}_{LS})^{T}\phi_{x_{i},a_{i}}\right)^{2}\leq 2\sum_{i=1}^{n}\left((\theta^{*})^{T}\phi_{x_{i},a_{i}}-(\widehat{\theta}_{LS})^{T}\phi_{x_{i},a_{i}}\right)\varepsilon_{i},$
where $\left\{\varepsilon_{i}\right\}_{i=1}^{n}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d. centered Gaussian distribution random variables. This is known as the basic inequality in regression. Denoting $\Phi\in\mathbb{R}^{n\times d}$ as the design data matrix and $y=(y_{1},\cdots,y_{n})^{T}$, $\varepsilon=(\varepsilon_{1},\cdots,\varepsilon_{n})^{T}$, we rewrite the model above as $y=\Phi\theta^{*}+\varepsilon$. Thus, what we have is equivalent to
$||\Phi\theta^{*}-\Phi\widehat{\theta}_{LS}||\leq 2\left\langle\varepsilon,v(\varepsilon)\right\rangle.$
where $v(\varepsilon)=\frac{\Phi\theta^{*}-\Phi\widehat{\theta}_{LS}}{||\Phi\theta^{*}-\Phi\widehat{\theta}_{LS}||}$. Let $U$ be a matrix with orthogonal columns such that its column space is the column space of $X$. Thus we get
$||\Phi\theta^{*}-\Phi\widehat{\theta}_{LS}||\leq 2\sup_{||a||\leq 1}\left\langle\varepsilon,Ua\right\rangle=2||U^{T}\varepsilon||_{2}.$
Because $\varepsilon$ is a Gaussian random vector and $U$ is orthogonal, so $U^{T}\varepsilon$ is another Gaussian random vector and $U^{T}\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{r})$ where $r=rank(X)$. Thus
$||\Phi\theta^{*}-\Phi\widehat{\theta}_{LS}||_{2}^{2}\leq 4||U^{T}\varepsilon||_{2}^{2}.$
Then we apply the Bernstein tail bound to achieve that with probability at least $1-\delta$,
$||U^{T}\varepsilon||_{2}^{2}\leq\sigma^{2}(r+2\sqrt{r\log(1/\delta)}+2\log(1/\delta))\leq\sigma^{2}(d+2\sqrt{d\log(1/\delta)}+2\log(1/\delta)).$
Thus we have
$\sum_{i=1}^{n}\mathsf{D}_{H}^{2}(f^{*}_{x_{i},a_{i}}||\widehat{f}^{n}_{x_{i},a_{i}})\leq\sigma^{2}(d+2\sqrt{d\log(1/\delta)}+2\log(1/\delta)).$
So we finish the proof. ∎
Proof of theorem 9.2.
For any function $f\in\mathcal{F}$, we define that
$Z_{i}=\frac{-1}{2}(\log f^{*}_{x_{i},a_{i}}(y_{i})-\log f_{x_{i},a_{i}}(y_{i})).$
Then we know that from lemma A.4, we know that with probability at least $1-\delta/|\mathcal{F}|$,
$\sum_{t=1}^{T}\frac{-1}{2}(\log f^{*}_{x_{i},a_{i}}(y_{i})-\log f_{x_{i},a_{i}}(y_{i}))\leq\sum_{t=1}^{T}\log\left(\mathbb{E}\left[\sqrt{\frac{f_{x_{i},a_{i}}(y_{i})}{f^{*}_{x_{i},a_{i}}(y_{i})}}\right]\right)+\log(|\mathcal{F}|/\delta).$
We further have that by the inequality $\log(1+x)\leq x$ that
$\log\left(\mathbb{E}\left[\sqrt{\frac{f_{x_{i},a_{i}}(y_{i})}{f^{*}_{x_{i},a_{i}}(y_{i})}}\right]\right)\leq\mathbb{E}\left[\sqrt{\frac{f_{x_{i},a_{i}}(y_{i})}{f^{*}_{x_{i},a_{i}}(y_{i})}}-1\right].$
On the other hand, we have
$\mathbb{E}\left[\sqrt{\frac{f_{x_{i},a_{i}}(y_{i})}{f^{*}_{x_{i},a_{i}}(y_{i})}}-1\right]\leq-\mathsf{D}_{H}^{2}(f_{x_{i},a_{i}}(y_{i})||f^{*}_{x_{i},a_{i}}(y_{i})).$
Plugging this back and by some algebra, we have
$\sum_{t=1}^{T}\mathsf{D}_{H}^{2}(f_{x_{t},a_{t}}(y_{t})||f^{*}_{x_{t},a_{t}}(y_{t}))\leq\sum_{t=1}^{T}\frac{1}{2}\left(\log(f^{*}_{x_{t},a_{t}}(y_{t}))-\log(\widehat{f}^{t}_{x_{t},a_{t}}(y_{t}))\right)+\log(|\mathcal{F}|/\delta)\leq\log(|\mathcal{F}|/\delta).$
The last inequality is by the definition of MLE. ∎